### v1

# Preference Robustness for DPO with Applications to Public Health

Abstract
We study an LLM fine-tuning task for designing reward functions for sequential resource allocation problems in public health, guided by human preferences expressed in natural language. This setting presents a challenging testbed for alignment due to complex and ambiguous objectives and limited data availability. We propose DPO-PRO, a robust fine-tuning algorithm based on Direct Preference Optimization (DPO), which accounts for uncertainty in the preference distribution using a lightweight Distributionally Robust Optimization (DRO) formulation. Unlike prior DRO-based DPO methods, DPO-PRO is significantly less conservative. We evaluate DPO-PRO on a real-world maternal mobile health program operated by the non-profit organization ARMMAN, as well as on standard alignment benchmarks. Experimental results demonstrate that our method consistently improves robustness to noisy preference signals compared to existing DPO variants. Moreover, DPO-PRO achieves comparable performance to prior self-reflection-based baseline for reward function design, while requiring significantly lower inference-time cost.
1 Introduction
Limited resource allocation is a core challenge in public health, where decision-makers must prioritize individuals for care or intervention under tight financial and operational constraints. In maternal healthcare, for instance, preventative care awareness programs have been shown to significantly reduce maternal mortality (Helpmum 2025; Jahan et al. 2018; ARMMAN 2025), yet are often implemented by non-profit organizations that must serve large populations with limited resources. To support such efforts, it is crucial to develop allocation policies that make effective sequential decisions based on individual-level attributes such as age, income, or health status.
Reinforcement learning (RL) offers a powerful framework for learning allocation policies in complex, dynamic environments. A widely used model in this context is the Restless Multi-Armed Bandit (RMAB) (Niño-Mora 2023), a class of Markov Decision Processes (MDPs) well-suited for sequential decision-making under resource constraints. However, in RL, it is well-recognized that designing an appropriate reward function is critical (Singh, Lewis, and Barto 2009; Booth et al. 2023). In public health applications, decision-makers’ priorities (e.g., which populations to prioritize) must be encoded into the reward function, so that resulting policy outcomes align with long-term public health goals. Crafting a reward function that accurately captures these priorities is often labor-intensive, demanding manual tuning, repeated simulation, and domain expertise. As public health goals shift over time, this design process must be repeated, limiting the scalability and adaptability of RL-based solutions in dynamic, real-world environments.
Figure 1: Overview of the implementation of DPO-PRO to public health applications such as the maternal health ARMMAN program in Section 5.2.
Recently, several works explored using large language models (LLMs) to automate reward function design for sequential decision-making problems (Mirchandani, Karamcheti, and Sadigh 2021; Goyal, Niekum, and Mooney 2019; Carta et al. 2022; Ma et al. 2024; Behari et al. 2024; Verma et al. 2025). In these works, an LLM interprets a user’s high-level objective (e.g., desired policy outcomes) in natural language and outputs a reward function that guides RL policy training. This enables efficient workflow for developing human-aligned RL policies. Recent studies demonstrated that LLMs can effectively generate reward functions for RMAB problems in public health domains (Behari et al. 2024; Verma et al. 2025).
Existing approaches often rely on inference-time optimization via self-reflection (Shinn et al. 2023), where the LLM iteratively refines its outputs based on simulated feedback. While this strategy offers flexibility, it incurs substantial inference-time costs due to repeated rounds of simulation and reasoning. Furthermore, in non-profit settings, data sensitivity constraints often limit the use of large proprietary models, leading organizations to adopt smaller open-source alternatives that may be less capable off-the-shelf but can be finetuned. Because self-reflection is highly dependent on model strength, its effectiveness may be limited in these contexts. Fine-tuning offers a more practical alternative by directly improving model capabilities while avoiding the high inference costs and computational overhead that become prohibitive in large-scale public health deployments (Wu et al. 2024), making it well-suited for resource-constrained domains such as public health (Zhao et al. 2025).
Thus, we aim to move beyond inference-time methods and instead fine-tune an LLM to more effectively translate natural language preferences into high-performing reward functions for RMABs. A promising approach is Direct Preference Optimization (DPO) (Rafailov et al. 2023), which enables LLM fine-tuning from pairwise preference annotations. Recent work has also leveraged LLM-as-a-judge to scale annotation as a proxy for human feedback (Zhang et al. 2025; Tan et al. 2024; Cui et al. 2024; Zhu et al. 2024). However, applying these techniques to the public health domain introduces significant challenges, particularly due to the presence of noise in the preference signal.
First, the objectives in the public health domain are often ambiguous or subjective (e.g., “prioritize older individuals”). Second, reward design in RMABs demands complex reasoning, as the quality of a proposed reward function is not evaluated in isolation, but rather by the policy it induces, increasing the complexity of assessing preference. Third, real-world public health datasets are typically smaller, increasing the risk of misalignment due to overfitting and reward hacking, which is a common problem with reinforcement learning for human feedback methods such as DPO (Gheshlaghi Azar et al. 2024; Amini, Vieira, and Cotterell 2024; Yang et al. 2024; Xiao et al. 2025). However, public health is a high-stakes domain, where poorly aligned policies can lead to serious real-world consequences.
Figure 2: DPO-PRO achieves comparable reward function performance to DLM on various ARMMAN maternal health tasks (Section 5.3) while providing (i) constant inference time scaling with only a finetuned LLama 8B LLM and (ii) removing external high-end LLM APIs for deployment.
To tackle these challenges, we introduce DPO-PRO, a lightweight and distributionally robust enhancement of DPO (Rafailov et al. 2023). Its application to public health settings is summarized in Figure 1. DPO-PRO incorporates uncertainty in the preference distribution using an efficient yet effective Distributionally Robust Optimization (DRO) formulation (Rahimian and Mehrotra 2022), ensuring that the fine-tuned policy remains reliable under imperfect preference annotation. Unlike prior DRO-based DPO methods (Wu et al. 2025; Xu et al. 2025; Mandal, Sasnauskas, and Radanovic 2025), our approach avoids excessive conservatism and only adds negligible computational overhead. We further show that the resulting robust loss can be interpreted as a regularized DPO loss that penalizes model overconfidence and weak preference signals.
We evaluate DPO-PRO in a simulated environment based on real-world maternal health application involving ARMMAN, a nonprofit organization that delivers preventative care messages via mobile health programs (ARMMAN 2025). Health workers can place a limited number of live service calls each week to boost beneficiary engagement. Prior work has modeled this problem using RMABs, where each beneficiary’s engagement evolves over time and the objective is to allocate $K$ live calls per week to maximize long-term engagement. Our experiments demonstrate that DPO-PRO is more robust to noisy preference supervision than both vanilla DPO and prior DRO-based approaches across a range of public health objectives. We also show that our method shows comparable performance with the Decision Language Model (DLM), a previous method based on self-reflection that was applied in the context of ARMMAN (Behari et al. 2024), but requires significantly lower inference-time cost as shown in Figure 2. Additionally, our finetuned LLM has the benefit of removing the need for external LLM APIs during deployment, which is required when working with sensitive or private data.
2 Related Works
A growing body of work has explored robustness in RLHF and DPO, particularly by leveraging DRO. Wu et al. (2025) introduce a DRO-based DPO formulation using a KL-divergence-based ambiguity set. Xu et al. (2025) extend this approach to both KL and Wasserstein distances, while Mandal, Sasnauskas, and Radanovic (2025) explore robustness under total variation distance, applying it to both DPO and standard two-stage RLHF.
In contrast to these DRO-based approaches, other works assume explicit corruption models. For example, Chowdhury, Kini, and Natarajan (2024) assume that preference labels are flipped according to a known noise rate, and Bukharin et al. (2024) model preferences using a Bradley–Terry (BT) framework (Bradley and Terry 1952) with added noise on the reward differences.
Hong et al. (2024) use a DRO framework to develop adaptive loss functions for reward model learning, but focus is not on distributional robustness. Zhan et al. (2024) introduce robustness by constructing confidence sets over learned reward functions, followed by pessimistic policy optimization.
Broadly, existing robustness techniques fall into two categories: (1) methods relying on strong assumptions about noise models, which may be unrealistic in practice, and (2) DRO-based methods that avoid explicit corruption modeling but hedge against worst-case distributions. As discussed in Section 3.2, standard DRO approaches tend to be overly conservative and often require solving challenging min–max optimization problems. For tractability, prior methods rely on heuristic approximations (Wu et al. 2025; Xu et al. 2025), potentially undermining theoretical guarantees or practical robustness. In contrast, we provide efficient approach DPO-PRO, with theoretical guarantees, and simultaneously illustrate its applicability in real world public health settings.
3 Preliminaries
3.1 Data Distribution
In DPO, each data point is a tuple $(x,y_{1},y_{2},c)$. A prompt $x\in\mathcal{X}$ is drawn from a distribution $\mu$. Given $x$, responses $y_{1},y_{2}\in\mathcal{Y}$ are sampled independently from a policy $\pi(\cdot|x)$. A label $c\in\{1,-1\}$ is a Bernoulli random variable indicating whether response $y_{1}$ is preferred over $y_{2}$, with preference probability given by $p^{*}(y_{1}\succ y_{2}|x)$. The distribution $p^{*}$ represents the ground-truth preference distribution that the user intends the language model to align with.
In summary, the data-generating process involves three components: (1) the prompt distribution, (2) the response distribution, and (3) the preference distribution. This yields the following joint distribution: $P(x,y_{1},y_{2},c)\;=\;\mu(x)\,\pi(y_{1}|x)\,\pi(y_{2}|x)\,\bigl{[}\mathbb{I}_{\{c=1\}}\,p^{*}(y_{1}\succ y_{2}\mid x)\;+\;\mathbb{I}_{\{c=-1\}}\,p^{*}(y_{2}\succ y_{1}\mid x)\bigr{]}.$ We often omit the dependence of $p^{*}$ (and other preference distributions) on $(x,y_{1},y_{2})$ and simply write $p^{*}$ instead of $p^{*}(y_{1}\succ y_{2}|x)$ when the meaning is clear from context.
3.2 (Distributionally Robust) DPO
In standard RLHF, the first step is to learn a reward function $R_{\phi}$ from a fixed dataset $\mathcal{D}$. This is typically done by minimizing the negative log-likelihood:
$\mathbb{E}_{(x,y_{1},y_{2},c)\sim\mathcal{D}}\left[\log\sigma\left(c\cdot\left(R_{\phi}(x,y_{1})-R_{\phi}(x,y_{2})\right)\right)\right],$
where $\sigma$ is the sigmoid function.
Once the reward model is trained, a policy $\pi_{\theta}$ is optimized to maximize the expected reward while remaining close to a reference policy $\pi_{ref}$, typically via Proximal Policy Optimization (PPO) (Schulman et al. 2017):
$\max_{\pi_{\theta}}\mathbb{E}_{(x,y)\sim\pi_{\theta}}\left[R_{\phi}(x,y)\right]-\beta\,\mathrm{KL}(\pi_{\theta}\,\|\,\pi_{\mathrm{ref}}),$
where $\beta$ is a regularization parameter controlling the strength of the KL penalty.
DPO simplifies this pipeline by combining the two stages into a single objective that directly optimizes the policy from pairwise preference data. The DPO loss $\mathcal{L}_{\text{DPO}}(\pi_{\theta})$ is given by:
$\mathbb{E}_{(x,y_{1},y_{2},c)\sim\mathcal{D}}\!\bigl{[}\log\sigma\!\bigl{(}c\,\beta\Bigl{(}\log\tfrac{\pi_{\theta}(y_{1}|x)}{\pi_{\mathrm{ref}}(y_{1}|x)}-\log\tfrac{\pi_{\theta}(y_{2}|x)}{\pi_{\mathrm{ref}}(y_{2}|x)}\Bigr{)}\bigr{)}\bigr{]}.$
For brevity, we define the per-sample loss as:
$\ell_{\theta}(x,y_{1},y_{2},c)=-\log\sigma\!\Bigl{(}c\,\beta\!\Bigl{(}\log\tfrac{\pi_{\theta}(y_{1}|x)}{\pi_{\mathrm{ref}}(y_{1}|x)}-\log\tfrac{\pi_{\theta}(y_{2}|x)}{\pi_{\mathrm{ref}}(y_{2}|x)}\Bigr{)}\Bigr{)}.$
Using this expression, $\mathcal{L}_{\text{DPO}}(\pi_{\theta})$ is also equivalent to
$\mathbb{E}_{(x,y_{1},y_{2})\sim\mathcal{D}}[p^{*}\ell_{\theta}(x,y_{1},y_{2},1)+(1-p^{*})\ell_{\theta}(x,y_{1},y_{2},-1)].$
When the context is clear that the triplet $(x,y_{1},y_{2})$ and the parameter $\theta$ is fixed, we also use $\ell_{1}$ and $\ell_{-1}$ to denote $\ell_{\theta}(x,y_{1},y_{2},1)$ and $\ell_{\theta}(x,y_{1},y_{2},-1)$, respectively.
As discussed earlier, DPO is vulnerable to overfitting and noise in the data. To address this, recent work has proposed distributionally robust objectives of the form $\max_{P\in\mathcal{Q}(\mathcal{D})}\mathcal{L}_{{P}}(\pi_{\theta}),$ where $\mathcal{Q}(\mathcal{D})$ denotes an ambiguity set centered around the empirical distribution (data) $\mathcal{D}$ and the loss $\mathcal{L}_{{P}}$ is the DPO loss evaluated under the perturbed distribution $P$ rather than the original data distribution (Xu et al. 2025; Wu et al. 2025).
Conservatism and Computational Costs
The above DRO formulation hedges against shifts in the entire joint distribution on $(x,y_{1},y_{2},c)$. This allows the adversary to assign weights to highly unlikely prompts or responses. The outer minimization must then optimize against losses in these practically irrelevant regions, which might make the update overly conservative and slow improvement on the data the model actually sees. From a computational standpoint, solving the resulting min–max problem can be intractable in practice, and existing methods often rely on approximations that deviate from the original DRO formulation.
3.3 Reward Function Design Tasks for RMABs
In our setting, the prompt $x$ represents a desired policy outcome specified by a human prompter, for example, Focus on the young mothers by age and also focus on those with low income. See Figure 1. The responses $y_{1}$ and $y_{2}$ correspond to candidate reward functions for RMABs (details of the RMAB formulation are provided in the Appendix). The role of the LLM-annotator is to judge which of the two reward functions, $y_{1}$ or $y_{2}$, better aligns with the objective expressed in the prompt.
A key challenge in this task is that the quality of a reward function cannot be assessed in isolation. Rather, it must be evaluated based on the policy it induces when optimized, i.e., how effectively the resulting policy fulfills the prompter’s (often ambiguous) intent. This requires the annotator to infer the downstream effects of each reward function. Given the complexity and ambiguity involved, the resulting preference signals are inevitably noisy and should not be treated as fully reliable.
4 DPO with Preference Robustness
In this section, we introduce DPO-PRO, a distributionally robust version of DPO that specifically targets distributional shifts or other forms of noise in the preference distribution.
4.1 Uncertainties in the Preference Distribution
In each data point $(x,y_{1},y_{2},c)$, the triplet $(x,y_{1},y_{2})$ represents observed content, which typically comes from a well-understood and controllable data collection process. In general, $x$ is drawn from a pool of candidate inputs, and the response pair $(y_{1},y_{2})$ is generated by the reference policy $\pi_{ref}$ (usually a supervised-fine tuned model) for every sampled prompt $x$. As more data is collected, the uncertainty in this part of the data is expected to diminish. Noise in the prompts and responses becomes even less of an issue under iterative DPO methods, which are increasingly adopted in practice (Xiong et al. 2024; Cen et al. 2025; Xie et al. 2025).
In contrast, we argue that the preference distribution $c\sim p^{*}$ is inherently noisy in practice, and collecting more data is unlikely to resolve the underlying uncertainty in human preference. Even when humans are directly annotating, various sources of noise remain, including human subjectivity, irrational or inconsistent behavior, and temporal variability (e.g., the same annotator providing different judgments at different times). This noise is persistent and cannot be easily addressed through better data curation or increased data volume. In other words, the ground-truth preference distribution $p^{*}$ is an idealized and inaccessible object in practice. As mentioned in the introduction, this challenge is particularly pronounced in public health domains, where objectives are often ambiguous or subjective (e.g., “prioritize older individuals”), and reward design demands complex reasoning about long-term policy consequences that are difficult to assess accurately. Moreover, real-world public health datasets are typically smaller, further exacerbating preference uncertainty and increasing the risk of misalignment. Therefore, we argue that uncertainty in the preference distribution should be the primary motivation for incorporating robustness into DPO.
We assume that for each prompt and response pair $(x,y_{1},y_{2})$, we have access to a (potentially noisy) preference distribution $q(y_{1}\succ y_{2}|x)$. We apply DRO adjustment, assuming worst-case deviation from $q$ within a pre-specified chi-squared divergence ball.
Formally, we define the DRO loss $\mathcal{L}_{\text{DRO}}(\pi_{\theta})$ as:
$\mathbb{E}_{(x,y_{1},y_{2})\sim\mathcal{D}}\max_{p\in Q(x,y_{1},y_{2},\rho)}\mathbb{E}_{c\sim p}\Bigl{[}\ell_{\theta}(x,y_{1},y_{2},c)\Bigr{]}$ (1)
where $Q(x,y_{1},y_{2}\rho)=\left\{p:\chi^{2}\big{(}p\,\|\,q(y_{1}\succ y_{2}\mid x)\big{)}\leq\rho\right\}$. Our formulation avoids unnecessary robustification over the joint distribution on $(x,y_{1},y_{2})$, and assumes that this part of the data is reliable. Furthermore, unlike standard DRO methods that can add significant computational burden (Rahimian and Mehrotra 2022), our formulation introduces negligible additional cost as shown next.
4.2 Computing Worst-case Distribution
Consider a single data point $(x,y_{1},y_{2})$. We assume access to an estimate of the probability that one response is preferred over the other: $q(y_{1}\succ y_{2}|x)\in(0,1)$.
Given this probability $q$, we define the following per-sample optimization problem:
$\max_{p\in[0,1]}$ $p\,\ell_{\theta}(x,y_{1},y_{2},1)+(1-p)\,\ell_{\theta}(x,y_{1},y_{2},-1)$ (2)
s.t. $\frac{(p-q)^{2}}{q(1-q)}\;\leq\;\rho$
This problem finds the worst-case distribution within the chi-squared ambiguity set centered at $q$, $\left\{p:\chi^{2}\big{(}p\,\|\,q(y_{1}\succ y_{2}\mid x)\big{)}\equiv\frac{(p-q)^{2}}{q(1-q)}\leq\rho\right\}$. Let $\hat{p}(y_{1}\succ y_{2}|x)$ denote the optimal solution to this problem. Using this, the robust loss $\mathcal{L}_{\text{DRO}}(\pi_{\theta})$ can be equivalently expressed as
$\mathbb{E}_{(x,y_{1},y_{2})\sim\mathcal{D}}\Bigl{[}\hat{p}\ell_{\theta}(x,y_{1},y_{2},1)+(1-\hat{p})\ell_{\theta}(x,y_{1},y_{2},-1)\Bigr{]}.$
Eq (2) is a one-dimensional optimization problem with a linear objective, which admits a simple closed-form solution:
$\hat{p}(y_{1}\succ y_{2}|x)=\begin{cases}\min\left\{1,q+\sqrt{\rho\,q(1-q)}\right\},\ \text{if }\ell_{1}\geq\ell_{-1},\\[6.0pt] \max\left\{0,q-\sqrt{\rho\,q(1-q)}\right\},\ \text{if }\ell_{1}<\ell_{-1}.\end{cases}$
Consider the case where $\ell_{1}\geq\ell_{-1}$, which implies that under the current policy $\theta$, the model assigns a higher likelihood to $y_{2}$ being preferred over $y_{1}$. To increase the loss, the adversary seeks to shift the preference probability $q(y_{1}\succ y_{2}\mid x)$ in the opposite direction. That is, it tries to increase $q$ to emphasize a preference for $y_{1}$, which contradicts the model’s belief. Consequently, the worst-case distribution becomes $q+\sqrt{\rho\,q(1-q)}$. Since probabilities must remain within the unit interval, we clip the value at 1, yielding the final expression $\min\left\{1,\;q+\sqrt{\rho\,q(1-q)}\right\}$.
4.3 Efficient Optimization of the Loss
The worst-case probability obtained from the optimization problem above, $\hat{p}$, replaces the preference probability in the per-sample DPO gradient as well:
$\hat{p}\,\nabla_{\theta}\ell_{\theta}(x,y_{1},y_{2},1)\;+\;(1-\hat{p})\,\nabla_{\theta}\ell_{\theta}(x,y_{1},y_{2},-1).$ (3)
Proposition 4.1.
Eq (3) provides an unbiased gradient estimate of the DRO loss in Eq (1).
Note that we do not formally differentiate through the inner maximization in Eq (1) with respect to $\theta$. However, Danskin’s theorem (Bertsekas 1999) justifies our approach. For a fixed $\theta$, we may solve the inner maximization and directly substitute the resulting worst-case $\hat{p}$ into the gradient expression.
We emphasize the resulting DRO gradient is both exact and computationally efficient. Unlike prior work, our approach does not introduce any approximation or heuristic in minimizing the DRO objective. Moreover, our distributional robustness is applied only to the preference distribution $q$, rather than the full data-generating distribution on $(x,y_{1},y_{2},c)$. Hence, the resulting method is significantly less conservative than earlier DRO-based DPO approaches, and matching more accurately the type of distribution shift we expect in the true human preferences in our public health application as seen later.
4.4 Analysis of the Loss Function
In this section, we analyze the robust loss $\mathcal{L}_{\text{DRO}}$ and show that it is equivalent to the original DPO loss augmented with a regularization term. This interpretation provides insight into how the DRO formulation affects model learning.
Proposition 4.2.
The DRO loss $\mathcal{L}_{\text{DRO}}$ is equivalent to regularizing the original DPO loss $\mathcal{L}_{\text{DPO}}$ as:
$\mathcal{L}_{\text{DRO}}=\begin{cases}\mathcal{L}_{\text{DPO}}+\min\left\{1-q,\;\sqrt{\rho\,q(1-q)}\right\}(\ell_{1}-\ell_{-1}),\\[4.0pt] \hskip 30.00005pt\text{if }\ell_{1}\geq\ell_{-1},\\[8.0pt] \mathcal{L}_{\text{DPO}}+\min\left\{q,\;\sqrt{\rho\,q(1-q)}\right\}(\ell_{-1}-\ell_{1}),\\[4.0pt] \hskip 30.00005pt\text{if }\ell_{1}<\ell_{-1}.\end{cases}$
We now analyze the behavior of the regularization term, assuming that $\ell_{1}\geq\ell_{-1}$. The opposite case follows symmetric logic.
(1) Uncertainty-Weighted Coefficient.
The first factor $\min\left\{1-q,\;\sqrt{\rho\,q(1-q)}\right\}$ penalizes cases where the preference signal is uncertain (i.e., when $q$ is close to 0.5). This term is largest when $q$ is near 0.5 and diminishes as $q\to 0$ or $q\to 1$. Its magnitude is modulated by the robustness radius $\rho$, with larger $\rho$ leading to a stronger penalty.
(2) Model Confidence.
The second factor $(\ell_{1}-\ell_{-1})$ reflects the log-odds of the current model’s preference for $y_{2}$ over $y_{1}$. A large value indicates strong model confidence. This term increases when the model becomes more certain about the preference, regardless of the ground-truth label.
(3) Combined Effect.
The regularization term penalizes the model for being overly confident in its preferences when the preference signal $q$ is ambiguous. In contrast, when the signal is strong (i.e., $q$ is close to 0 or 1), the penalty diminishes, allowing the model to express stronger preferences. In effect, the DRO loss encourages calibrated learning: the model is allowed to be confident only when the preference signal is also confident.
4.5 Obtaining the Soft Score $q$ in Practice
The DRO adjustment described above requires a soft score $q(y_{1}\succ y_{2}|x)$, rather than a binary annotation for each prompt and response pair $(x,y_{1},y_{2})$. In this work we focus on settings where preferences are evaluated by LLM judges. Soft scores can be obtained from LLM judges in several ways: (1) producing multiple binary preference judgments through repeated queries and averaging the outcomes; (2) directly outputting a numerical preference score via appropriate prompting; or (3) extracting the log-probabilities assigned to tokens representing the preferences (e.g., 1 for $y_{1}\succ y_{2}$ and -1 for $y_{2}\succ y_{1}$), and applying a softmax transformation to estimate a preference distribution (Lee et al. 2024). We provide further discussion on annotation formats and the connection to robustness in the Appendix.
5 Experiments
We validate DPO-PRO in two different settings. First, we validate the methodology on a standardized benchmark for preference-based LLM finetuning. Next, we adapt this experiment for the ARMMAN maternal health dataset (ARMMAN 2025) in the following section.
5.1 Benchmarks on General Alignment Datasets
We use the recent UltraFeedback alignment benchmark dataset (Cui et al. 2024), consisting of 60,000 top-quality human pairwise preference annotations that cover a wide variety of tasks and instruction types, rendering it a broad benchmark for examining alignment research.
Noise in the training data
Our primary objective is to evaluate how effectively preference-based LLM fine-tuning methods handle shifts in preference distributions between training and evaluation data. We use $q^{*}$ denote the true preference distribution used at test time. For our experiments, we define $q^{*}$ using the state-of-the-art preference reward model Eurus 7B(Yuan et al. 2025), which has been trained for the UltraFeedback dataset (Cui et al. 2024). We compute $q^{*}$ using the BT model.
To simulate noise during training, we construct a noisy preference distribution $q_{\alpha}$ for training as the mixture
$q_{\alpha}:=q^{*}(1-\alpha)+(1-q^{*})\alpha.$ (4)
Intuitively, $q_{\alpha}$ “flips” the preference distribution with probability $\alpha$. We refer to this type of noise interchangeably as label switching. During training, we have access to $q_{\alpha}$ but not to $q^{*}$, with $\alpha$ being unknown. We include additional experiments with alternative forms of noise in Appendix.
Training details
We base our experiments on the recent Phi 3-mini 3B model without instruction (Abdin et al. 2024). We initially train for two epochs of supervised fine tuning (SFT), followed by one epoch of DPO, with the baselines discussed below. All experiments are performed on a Nvidia A100 GPU 40 GB with batch size 2. Refer to the technical appendix for full details on the training procedure such as weight decay, learning rate scheduler, low-rank adaptation parameters, quantization, etc and experiments conducted on the Llama 3 8B model (Grattafiori et al. 2024), not shown here due to space limitations.
Baselines
We compare DPO-PRO against vanilla DPO (Rafailov et al. 2023) and the two robust baselines: rDPO (Chowdhury, Kini, and Natarajan 2024) and DrDPO (Wu et al. 2025). We implement DPO-PRO with a value of $\rho=0.1$ for the $\chi^{2}$-divergence. We train for two noise levels ($\alpha=0.1$ and $\alpha=0.3$), termed low and high respectively, and also for the noiseless case $\alpha=0$.
Evaluation metrics
After the DPO step, we generate a response for every prompt $x$ in the evaluation dataset $y_{g}\sim\pi_{\theta}(x)$. Then, we compute the reward $R_{\phi}(x,y_{g})$. We focus on two evaluation metrics, the win rate, which is computed by comparing the generated prompt with the chosen response $y_{c}$ in the dataset as $\texttt{win\_rate}=\frac{1}{N_{\mathrm{eval}}}\sum_{i=1}^{N_{\mathrm{eval}}}\mathbb{I}(R_{\phi}(x,y_{g})>R_{\phi}(x,y_{c})),$ where $\mathbb{I}$ is the indicator function. The next metric is the generation evaluation reward, computed simply as $\texttt{eval\_reward}=\frac{1}{N_{\mathrm{eval}}}\sum_{i=1}^{N_{\mathrm{eval}}}R_{\phi}(x,y_{g})$.
Results
Table 1(b) presents the results. Across all noise levels, DPO-PRO consistently achieves highest performance in both win rate and evaluation reward. Notably, it outperforms all baselines even in noise-free setting, highlighting robustness without sacrificing standard performance.
We found it surprising that performance of all methods increased with additional noise. This can be explained by the fact that DPO in fixed (off-policy) data is known to be highly sensitive to reward hacking (Gheshlaghi Azar et al. 2024; Amini, Vieira, and Cotterell 2024; Yang et al. 2024; Xiao et al. 2025). An analogous result for DPO-based methods can be found in Wu et al. (2025); Bukharin et al. (2024). The noise seemingly acts as a regularization, preventing model from overfitting to spurious patterns in training data and instead learning more generalizable features. This phenomenon is consistent with recent theoretical work showing controlled noise injection can improve generalization by encouraging simpler decision boundaries (Li et al. 2025).
Table 1: Results on UltraFeedback under label-flip noise. Experiments run with Phi-3-mini. Mean and standard error over 2000 evaluation samples are shown.
(a) Win rate ($\uparrow$)
Method No noise Low noise (0.1) High noise (0.3)
[…]
(b) Generation Evaluation Reward ($\uparrow$)
Method No noise Low noise (0.1) High noise (0.3)
[…]
5.2 Benchmarks on Maternal Health Data
We apply our proposed method in the context of mobile health program run by ARMMAN (2025). Specifically, we have anonymized beneficiary listenership data from a quality-improvement study conducted by ARMMAN [^1] The authors have been access to this restricted dataset under a data usage agreement.. The data consists of registration information of beneficiaries (mothers) containing their socio-demographic features of age, income education as well as their preferences on call slot times and language. Additionally, we have information on beneficiaries’ interaction with automated voice calls and live service calls made by health-workers. We model the problem of allocating live service calls as a RMAB problem. For all our experiments, we consider a sample of population with $2100$ beneficiaries and $K=210$ as the budget, representing the number of live service calls that can be made every week. Exact policy optimization for RMABs is generally known to be intractable (Papadimitriou and Tsitsiklis 1987). However, Whittle index policy (Whittle 1988) is a popular heuristic to obtain an asymptotically optimal policy. Please refer also to the technical appendix for additional details.
Preference Dataset Construction
We generate a preference dataset of health worker prioritization commands, preferred, and rejected reward functions as shown in Figure 1. Since it is costly to generate this dataset through human annotation, we use an LLM judge with ChatGPT 4o-mini (OpenAI 2024) as follows: i) query LLM-judge to obtain 20 candidate reward functions that align with the prioritization command, ii) for each candidate reward function, solve RMAB problem using Whittle index method and generate trajectory outputs, iii) sample 50 pairs of reward functions from this set and query the LLM-judge to select preferred reward function, iv) perform this query 10 times to estimate the LLM’s uncertainty over preferences. Finally, we obtain a dataset of 9500 preferred and rejected reward function responses over 190 prioritization commands (see Appendix for more details).
Noise injection, training details, and baselines
We adopt the same noise injection scheme to obtain $q_{\alpha}$ as described in Section 5.1 and Eq (4). We fine-tune a Llama 3 (8B) model, keeping the learning rate, DPO reward temperature, $\chi^{2}$-divergence radius, and other hyperparameters identical to those used in the UltraFeedback benchmark experiments. Given poor performance of rDPO in Table 1(b), we omit this baseline in this experiment.
Evaluation
To evaluate the quality of a generated reward function, we follow a procedure analogous to the dataset creation: i) solve the resulting RMAB problem using the Whittle Index method to generate policy trajectories, ii) use LLM judge (GPT-4o-mini) to evaluate these policy outcomes. Specifically, we compute the win rate by comparing the policy induced by our generated reward function to the policy obtained from the reward function in the training data.
Table 2: Examples of reward functions produced by Vanilla-DPO and DPO-PRO when trained under noise. DPO-PRO produces more conservative preference interpretations.
Task 1: Prefer both young and elderly beneficiaries
Vanilla-DPO s + 3 * (youngest_age or second_youngest_age or oldest_age) +
    2 * (lowest_education or second_lowest_education or third_lowest_education) +
    (lowest_income or second_lowest_income or third_lowest_income)
DPO-PRO s + 3 * (youngest_age or second_youngest_age or oldest_age) +
    2 * (lowest_education or second_lowest_education or third_lowest_education)
Task 2: Prioritize midday calls and NGO-registered users
Vanilla-DPO s + 3 * (10_30-12_30pm and NGO_registered) +  2 * (12_30-3pm and NGO_registered)
DPO-PRO s + 3 * (12_30-3pm and NGO_registered)
Figure 3: Winrates for RMAB environment under label flipping noise. We present more results in the Appendix.
Results
Table 3 shows the results. In zero-noise setting, vanilla-DPO outperforms robust methods. However, as noise level increases, DPO-PRO outperforms vanilla DPO and DrDPO. We also perform a task-level analysis of winrates in Figure 3. We note that across all tasks, under the presence of noise, robust DRO methods are better than vanilla-DPO. Especially, in 5 out of 8 evaluation task categories, DPO-PRO outperforms baselines under the presence of noise. This signifies the robustness of our approach.
To further illustrate the differences between vanilla DPO and DPO-PRO, we present response outputs from both models on two representative tasks from the evaluation dataset (Table 2). Many tasks contain attributes that can be interpreted in multiple ways. For instance, the concept of “young beneficiaries” can be identified directly through the age feature, but also indirectly via lower education level or income. Similarly, a task that expresses a preference for “midday calls” could refer explicitly to the 12:30–3:00 PM time slot, but might also include the 10:30–12:30 PM window, depending on interpretation.
In such cases, we observe that vanilla-DPO favors uncommon interpretations of ambiguous attributes based on spurious correlations in noisy preference data. In contrast, DPO-PRO adopts a cautious strategy by selecting narrower definition, aiming to be robust to uncertainty in the preference signal at evaluation.
5.3 Comparison with Decision-Language Model
Table 4 compares DPO-PRO with the Decision Language Model (DLM), a self-reflection-based approach introduced in (Behari et al. 2024), using two evaluation metrics: (1) win rate, determined by an external LLM judge (GPT-4o-mini), and (2) reward, computed using ground-truth human-designed reward functions. While DLM slightly outperforms DPO-PRO on both metrics, the results indicate that DPO-PRO achieves comparable alignment quality despite its simplicity and significantly lower inference-time cost.
Figure 2 provides detailed comparison of inference-time cost between DPO-PRO and DLM. Unlike DLM, whose inference cost scales linearly with number of arms due to repeated simulation and self-reflection, DPO-PRO incurs a constant inference cost regardless of problem size. This efficiency gap becomes increasingly significant in large-scale settings, making DPO-PRO more scalable. While overall performance is comparable or mixed across evaluation metrics, minimal inference-time overhead makes DPO-PRO especially well-suited for real-world deployment.
Table 3: Winrate (%, $\uparrow$) comparision for ARMMAN environment for different methods under no noise and high noise settings, aggregated across all tasks.
Method Vanilla-DPO DrDPO DPO-PRO
No Noise $\bf 55\pm 2.3$ $40\pm 1.85$ $35.3\pm 1.83$
High Noise (0.3) $45\pm 2.03$ $54.8\pm 1.84$ $\bf 63.9\pm 1.94$
Table 4: Comparison of Winrate and Reward
Method Winrate Ground Truth Reward
DLM $35.5\pm 5.02$ $6073.04\pm 88.5$
DPO-PRO $32.2\pm 4.92$ $5687.3\pm 112.3$
6 Conclusion
Motivated by reward function design tasks in RMABs for public health, we introduced DPO-PRO, a robust fine-tuning algorithm that enhances DPO by explicitly modeling uncertainty in preference distributions through a lightweight DRO formulation. Our method is particularly well-suited for high-stakes decision-making settings like public health, where preference annotations are inherently noisy and often limited. Compared to existing DRO-based approaches, DPO-PRO avoids excessive conservatism while providing strong theoretical justification as a regularized variant of DPO. Through extensive evaluation on a real-world maternal health program and standard alignment benchmarks, we demonstrated that DPO-PRO improves robustness, and achieves competitive performance to inference-time methods such as self-reflection at a fraction of the computational cost. Additionally, contrary to inference-time methods, DPO-PRO removes the dependency on external LLM APIs at deployment which is essential when working with sensitive or private data. These results highlight the potential of DPO-PRO as a scalable and reliable tool for learning aligned reward functions in real-world policy learning tasks for public health.
Appendix A Codebase
We release the codebase for all our experiments on finetuning LLMs here: https://anonymous.4open.science/r/DPO-Pro-A33D/README.md
Appendix B Extensions and Generalizations
B.1 Beyond the Chi-squared Divergence
We use the chi-squared divergence to define the ambiguity set due to the simplicity of its resulting closed-form solution. However, alternative divergence measures can also be used and are likely to offer comparable computational efficiency as ours. This is because the per-sample optimization problem in Eq (2) remains a one-dimensional problem with linear objective. For example, using the KL divergence leads to a constraint of the form: $p\log\frac{p}{q}+(1-p)\log\frac{1-p}{1-q}\;\leq\;\rho$. In this case, the worst-case $\hat{p}$ can be computed by solving the equation $p\log\frac{p}{q}+(1-p)\log\frac{1-p}{1-q}\;=\;\rho$, which can be efficiently handled using a simple one-dimensional root-finding method.
B.2 Obtaining the Soft-score $q$
In many practical applications beyond ours, only a single binary annotation might be available per $(x,y_{1},y_{2})$. In this case, we effectively observe only one sample from the underlying Bernoulli distribution $q(y_{1}\succ y_{2}|x)$, which makes estimating $q$ challenging.
However, our method naturally accommodates binary annotations, which can be especially useful when part of the dataset is labeled by humans. For instance, a healthcare worker involved in the annotation process may choose to directly express their preferences, in which case requesting a binary label is more appropriate. We can slightly relax the definition of chi-squared divergence by rewriting the constraint as $(p-q)^{2}\leq\rho\,q(1-q),$ which reduces to $\hat{p}=q$ when $q\in\{0,1\}$. As a result, when hard binary labels are used, the DRO loss reduces to the standard DPO loss, introducing no additional regularization. This effectively means that we regularize only on LLM-generated annotations, while treating human-provided labels as ground truth.
To inject robustness in such cases, we can smooth the binary label by replacing the label $1$ with $1-\epsilon$, and 0 with $\epsilon$, where $\epsilon$ is a small hyperparameter. The DRO formulation can then be applied to this smoothed label. This makes the uncertainty-weighted coefficient uniform across all pairs (since $q$ only takes values $1-\epsilon$ or $\epsilon$), but the $(\ell_{1}-\ell_{-1})$ term still encourages the model to reduce overconfidence in its current preference.
Our method also applies to the case where off-the-shelf reward models are used for DPO (Guo et al. 2024; Ding et al. 2024; Lee et al. 2025; Gupta et al. 2025). In this setting, acquiring soft score $q(y_{1}\succ y_{2}|x)$ is straightforward. The scalar outputs of the reward model can be converted into a pairwise preference probability using the BT model. However, these preference probabilities can be noisy or uncalibrated. We discuss alternative annotation formats and how our approach adapts robustness to each setting in the Appendix.
B.3 Uncertainty-Weighted Coefficient
We visualize this uncertainty-weighted coefficient for various $\rho$ values in Figure 4.
Figure 4: Visualization of the uncertainty-weighted coefficient for various values of $\rho$. This coefficient attains its maximum near $q=0.5$ and decreases as $q\to 0$ or $q\to 1$. For small values of $\rho$, the maximum is achieved exactly at $q=0.5$, where $\sqrt{\rho\,q(1-q)}$ peaks. However, for larger $\rho$, the maximum occurs at the intersection point between the curves 1q1-q and $\sqrt{\rho\,q(1-q)}$. This is because $q+\sqrt{\rho\,q(1-q)}$ (the worst-case distribution under the chi-squared divergence constraint) may exceed 1 when $\rho$ is large, causing the worst-case distribution $\hat{p}$ to be clipped at 1. In such cases, the adversary’s perturbation $\hat{p}-q$ can be greater when $q<0.5$, leading the penalty term to peak at a value of $q$ smaller than 0.5.
Appendix C Omitted Proofs
C.1 Proof of Proposition 4.1
The result follows from Danskin’s theorem (Bertsekas 1999). To apply the theorem, we must verify that the inner maximization problem in Eq. (2) has a compact feasible set and admits a unique optimal solution.
To establish compactness, we need to show that the feasible set is both bounded and closed. Boundedness is immediate since $p\in[0,1]$. Closedness follows from the fact that the constraint function $p\mapsto(p-q)^{2}-\rho q(1-q)$ is continous in $p$. Hence, the set $(p-q)^{2}\leq\rho q(1-q)$ is closed. Since $p\in[0,1]$ is also closed and the intersection of closed sets is closed, the feasible set is closed. Uniqueness of the optimal solution $\hat{p}$ is also straightforward, since the objective is one-dimensional linear function in $p$.
C.2 Proof of Proposition 4.2
We begin by assuming the case $\ell_{1}\geq\ell_{-1}$. Then,
$\mathcal{L}_{\text{DRO}}$ $=\hat{p}\,\ell_{1}+(1-\hat{p})\,\ell_{-1}$
$=q\,\ell_{1}+(1-q)\,\ell_{-1}+(\hat{p}-q)(\ell_{1}-\ell_{-1})$
$=\mathcal{L}_{\text{DPO}}+\underbrace{\min\left\{1-q,\;\sqrt{\rho\,q(1-q)}\right\}}_{\text{Uncertainty-Weighted Coefficient}}(\ell_{1}-\ell_{-1}).$
For the opposite case, when $\ell_{-1}>\ell_{1}$, the derivation proceeds symmetrically.
$\mathcal{L}_{\text{DRO}}$ $=\hat{p}\,\ell_{1}+(1-\hat{p})\,\ell_{-1}$
$=\mathcal{L}_{\text{DPO}}+(\hat{p}-q)(\ell_{1}-\ell_{-1})$
$=\mathcal{L}_{\text{DPO}}-\min\left\{q,\,\sqrt{\rho\,q(1-q)}\right\}(\ell_{1}-\ell_{-1})$
$=\mathcal{L}_{\text{DPO}}+\min\left\{q,\,\sqrt{\rho\,q(1-q)}\right\}(\ell_{-1}-\ell_{1}).$
Appendix D Multi-armed Restless Bandits
A Multi-Armed Restless Bandit problem is characterized by N Markov Decision processes, each defined by the tuple $\{\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{T},\gamma\}$. Here, $\mathcal{S}$ represents the set of states, $\mathcal{A}$ represents the set of actions which are binary in $\{0,1\}$, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ represents the transition probability from current state, under an action to the next state. $\mathcal{R}:\mathcal{S}\to\mathbb{R}$ represents a reward function mapping state to a real number, and $\gamma$ represents the discounting factor. We also have a budget $K$ which represents the number of arms that can be chosen at every timestep. The objective of the Multi-Armed Restless Bandit problem then is to decide a vector of actions $\textbf{A}^{t}$ representing actions $a_{i}^{t}\in\mathcal{A}$ for every arm i in time step $t$ such that we maximize long term reward under the budget constraint:
$J(s)=\max_{\textbf{A}\in\mathcal{A}}\left(\sum_{i=1}^{N}R_{i}(s_{i})+\gamma\mathbb{E}[J(s^{\prime})\mid s,\textbf{A}]\right)$
$s.t.\sum_{i=1}^{i=N}A_{i}^{t}\leq K\;\forall\;t=1\;\text{to}\;T$
where $s_{t}\sim T(s_{t-1},a_{t},.)$
Appendix E Experimental Details
E.1 Additional Comparison with Baselines
rDPO (Chowdhury, Kini, and Natarajan 2024) does not use a DRO framework. Instead, it assumes that the observed preference labels are flipped at a known rate $\epsilon\in[0,0.5)$. Under this assumption, they derive an unbiased estimator of the DPO loss by reweighting the terms as follows:
$\frac{1}{1-2\epsilon}\left[(1-\epsilon)\ell(x,y_{1},y_{2}1)-\epsilon\ell(x,y_{1},y_{2},-1)\right],$
where $y_{1}$ is the chosen response in the noisy data.
DrDPO applies DRO over the empirical data distribution using a KL-divergence-based ambiguity set:
$\min_{\pi_{\theta}}\max_{P\in\mathcal{Q}(\mathcal{D})}\mathbb{E}_{(x,y_{1},y_{2},c)\sim P}[\mathcal{L}_{\text{DPO}}(\pi_{\theta})],$ (5)
where $\mathcal{Q}(\mathcal{D}):=\left\{P:D_{\text{KL}}(P\|\mathcal{D})\leq\rho\right\}$. Using duality, the inner maximum (the objective to minimize) is approximated by:
$\beta^{\prime}\log\mathbb{E}_{(x,y_{1},y_{2},c)\sim\mathcal{D}}\left[\exp\left(\frac{1}{\beta^{\prime}}\mathcal{L}_{\text{DPO}}(\pi_{\theta})\right)\right],$ (6)
where $\beta^{\prime}$ is a hyperparameter.
E.2 Whittle Index Policy
Whittle index is a heuristic that is asymptotically optimal for planning in Multi-Armed Restless Bandits. To calculate whittle index, first we define the value function as
$V_{i}^{\lambda}(s)=\max_{a\in\{0,1\}}Q_{i}(s,a_{i},\lambda).$ (7)
Here, $Q_{i}(s,a_{i},\lambda)$ measures the expected discounted cumulative future reward where a reward compensation $\lambda$ is added to the reward when the passive action is taken. The Whittle index associated to the state $s_{i}$ is then defined as:
$W_{i}(s_{i}):=\inf_{m}\left\{Q_{i}(s_{i},a_{i}=0,m)=Q_{i}(s_{i},a_{i}=1,m)\right\}.$
The Whittle Index intuitively captures the value of taking an active action on an arm. We compute the Whittle Index using a binary search algorithm proposed by Qian et al. 2016.
Finally, as a policy, we choose at every timestep the beneficiaires with top-K whittle indices.
E.3 ARMMAN Real World Data Attributes
To conduct our experiments in the ARMMAN real-world setting, we use a dataset of 2100 beneficiaires as a subset from a quality improvement study conducted in January 2022. Specifically, we use transition probabilities of beneficiaries estimated from historical data and use these probabilities to simulate beneficiary behaviour. Additionally, for every beneficiaries, we have information in their sociodemographic attributes, registration information, as well as preference on time slots of calls and on the language in which they want to receive calls. Table 5 shows the beneficiary attributes we use as the set of features that LLM can use in the reward function.
Table 5: Features describing beneficiaries in the ARMMAN dataset
Feature Name Data Type
Enrollment gestational age Int
Enrollment delivery status Int
Gravidity (number of pregnancies) Int
Parity (number of viable pregnancies) Int
Live births count Int
Days to the first call Int
Ages 10–20 Binary
Ages 21–30 Binary
Ages 31–40 Binary
Ages 41–50 Binary
Ages 51–60 Binary
Speaks Hindi Binary
Speaks Marathi Binary
Speaks Gujurati Binary
Speaks Kannada Binary
Education level 1/7 – illiterate Binary
Education level 2/7 – 1–5th Grade Completed Binary
Education level 3/7 – 6–9th Grade Completed Binary
Education level 4/7 – 10th Grade Passed Binary
Education level 5/7 – 12th Grade Passed Binary
Education level 6/7 – Graduate Binary
Education level 7/7 – Post graduate Binary
Phone owner 0 (e.g., woman) Binary
Phone owner 1 (e.g., husband) Binary
Phone owner 2 (e.g., family) Binary
To be called from 8:30am–10:30am Binary
To be called from 10:30am–12:30pm Binary
To be called from 12:30pm–3:30pm Binary
To be called from 3:30pm–5:30pm Binary
To be called from 5:30pm–7:30pm Binary
To be called from 7:30pm–9:30pm Binary
NGO Binary
ARMMAN Binary
PHC Binary
Income bracket -1 (no income) Binary
Income bracket 1 (0–5000) Binary
Income bracket 2 (5001–10000) Binary
Income bracket 3 (10001–15000) Binary
Income bracket 4 (15001–20000) Binary
Income bracket 5 (20001–25000) Binary
Income bracket 6 (25001–30000) Binary
Income bracket 7 (30000–999999) Binary
E.4 RMAB Preference Data Generation
To generate the preference dataset for fine-tuning an LLM, we follow these steps:
  1. 1.
    We consider a set of 190 human preference commands similar to the ones proposed in Behari et al. 2024. Table shows a smaller sample of those commands.
  2. 2.
    For each of these commands, we use the DLM pipeline to generate a set of 20 relevant reward functions using prompt in Figure 6.
  3. 3.
    We sample a set of 50 pairs from these reward functions. For each of these pairs
    1. (a)
      solve the RMAB problem and generate trajectories for every reward function
    2. (b)
      ask an LLM Judge to decide the better reward function using the prompt in Figure 7.
  4. 4.
    We perform this query 10 times to estimate LLM’s uncertainty in preference for every pair of chosen and rejected reward functions.
E.5 Hyperparameters
Table 6 shows the hyperparameters used during training.
Table 6: Hyperparameter Configuration
Description RMAB UltraFeedback
Model Configuration
Maximum sequence length 2048 1536
LoRA rank parameter 64 64
LoRA scaling parameter 16 16
LoRA dropout rate 0.05 0.05
Data type bf16 bf16
SFT Training
Training epochs 2 2
learning rate 1e-5 1e-6
per device train batch size 8 4
number of devices 4 4
gradient accumulation steps 1 2
optimizer adamw_torch adamw_torch
weight decay 0.01 0.01
optimizer warmup steps 100 100
learning rate scheduler cosine cosine
DPO Training
training epochs 1 1
learning rate 1e-5 1e-6
per device train batch size 4 2
gradient accumulation steps 4 4
optimizer warmup steps 100 100
beta parameter 0.25 0.25
optimizer adamw_torch adamw_torch
warmup steps 100 100
learning rate scheduler cosine w/min_lr=1e-6 cosine w/min_lr=1e-7
Appendix F Additional Results and Ablations
Figure 5: RMAB environment win rates under adversarial label noise
F.1 Adversarial Noise
In the main text we defined and presented results for label flipping noise. Here we consider an additional noise distribution defined as
[…]
$U$ $\sim\mathrm{Unif}(0,\alpha)$
Similar to label flipping, adversarial noise tends to flip the labels, however, the mechanism is different as the distribution shift occurs as continuous change with an $\alpha$-ball of $q$.
The results of implementing DPO-PRO and the baselines with adversarial noise in the UltraFeedback dataset are shown in Table 7(b), consistent with the trends we observed with label flipping where additional noise improves training in general, and DPO-PRO being the best performer. Figure 5 shows the results for the public health (RMAB) application. We show the former results in a table and the latter in a table for consistency with the main text.
Table 7: Results on UltraFeedback under label-flip noise. Experiments run with Phi-3-mini. Mean and standard error over 2000 evaluation samples are shown.
(a) Win rate ($\uparrow$)
Method No noise Low noise (0.1) High noise (0.3)
[…]
(b) Generation Evaluation Reward ($\uparrow$)
Method No noise Low noise (0.1) High noise (0.3)
[…]
F.2 Additional Ablations
Effect of no Robustness ($\rho=0$)
Table 8 presents results for a variant of DPO-PRO with $\rho=0$, which effectively removes the DRO component. In this case, the method reduces to a form of DPO that uses soft preference probabilities $q$ instead of binary labels. Interestingly, the performance remains slightly better than vanilla DPO. We note that when $\rho=0$, the loss becomes a standard cross-entropy between probability scores.
Table 8: Flip Noise - Win Rate (%) and Evaluation Reward Comparison. DPO-PRO with $\rho=0$.
Method No noise $\uparrow$ Low noise (0.1) $\uparrow$ High noise (0.3) $\uparrow$
Win Rate (%) Eval Reward Win Rate (%) Eval Reward Win Rate (%) Eval Reward
[…]
UltraFeedback results on Llama3 8B
Table 9(b) shows the results of fitting the LLama 3 8B model instead of the Phi3-mini 3B model. As expected the results are slightly higher considering the larger and more capable initial neural network. However, the memory and computation requirements also increase.
Table 9: Results on UltraFeedback under label-flip noise run with Llama3 8B. Mean and standard error over 2000 evaluation samples are shown.
(a) Win rate ($\uparrow$)
Method No noise Low noise (0.1) High noise (0.3)
[…]
(b) Generation Evaluation Reward ($\uparrow$)
Method No noise Low noise (0.1) High noise (0.3)
[…]
Appendix G Prompts
table]box:rwgeneratorLLM prompt to generate reward function Prompt
Create a Python reward function for RL in phone call resource allocation to mothers in India, with the objective of prioritizing higher states and: {goal_prompt}.
The function should use ’state’ (value is either 0 or 1) and features ’agent_feats’ to direct the RL agent.
Here is a description of the features you may use along with the index in the ’agent_feats’ array: {PROMPT_FEAT_DESC_DICT}
Your task:
1. Write a simple, single-line Python reward function. - Exclude the word ’return’. - Exclude non-standard libraries. - Format your code with []: [YOUR FUNCTION].
2. Note that HIGHER states are always preferred, so ensure the reward increases as ’state’ increases.
3. Make sure the reward is always positive and increasing with state.
4. Avoid using bitwise operators. Use ’and’, ’or’ instead.
Here is an example to guide you: PROMPT_REW_EXAMPLE_DICT
Come up with a unique new reward for the specified goal: goal_prompt.
Figure 6: Prompt passed to the LLM to generate a reward function based on the context of problem scenario in Real World Domain.
table]box:rwgeneratorLLM Judge to Decide Preferred Response Prompt
My goal was to create a Python reward function for RL in resource allocation, with the objective of: Focus on those with high education. I tried several reward functions for this task. Below, I have the given reward function, and the corresponding distribution of reward achieved.
Below are the reward functions I used and their corresponding reward distributions: Function Number 0: Reward Function: -agent_feats[5] -agent_feats[6]-agent_feats[7]-agent_feats[8]-agent_feats[9]-agent_feats[10]-agent_feats[11]
Reflection: ’
Category: Age Ages 10-20: 121.73
Ages 21-30: 421.04
Ages 31-40: 244.49
Ages 41-50: 64.11
Ages 51-60: 10.58
Category: Income Income bracket 1 (e.g., 0-5000): 126.82
Income bracket 2 (e.g., 5001-10000): 373.62
Income bracket 3 (e.g., 10001-15000): 234.87
Income bracket 4 (e.g., 15001-20000): 77.40
Income bracket 5 (e.g., 20001-25000): 35.58
Income bracket 6 (e.g., 25001-30000): 2.58
Income bracket 7 (e.g., 30000-999999): 11.09
Category: Education Illiterate: 39.91
1-5th Grade Completed: 157.84
6-9th Grade Completed: 281.36
10th Grade Passed: 197.64
12th Grade Passed: 103.18
Graduate: 21.13
Post graduate: 60.89
Function Number 1: Reward Function: state * agent_feats[10]
Reflection: ’ Category: Age Ages 10-20: 134.22
Ages 21-30: 469.16
Ages 31-40: 270.44
Ages 41-50: 72.80
Ages 51-60: 11.96
Category: Income Income bracket 1 (e.g., 0-5000): 138.40
Income bracket 2 (e.g., 5001-10000): 414.44
Income bracket 3 (e.g., 10001-15000): 266.44
Income bracket 4 (e.g., 15001-20000): 85.33
Income bracket 5 (e.g., 20001-25000): 40.20
Income bracket 6 (e.g., 25001-30000): 2.80
Income bracket 7 (e.g., 30000-999999): 10.96
Category: Education Illiterate: 45.07
1-5th Grade Completed: 173.82
6-9th Grade Completed: 314.07
10th Grade Passed: 217.31
12th Grade Passed: 113.02
Graduate: 29.36
Post graduate: 65.93
Based on the above reward distributions and the given goal: Focus on those with high education., please identify the FUNCTION NUMBER of the most effective reward function. Provide your answer EXACTLY IN the following format: ’The best reward function is at number: [FUNCTION NUMBER]’.
Output:
The best reward function is at number: 1
Figure 7: Prompt passed to the LLM to choose a reward function based on the context of problem scenario in Real World Domain, the generated reward functions and the reward distribution corresponding to every reward function.

### v2

@@
 We propose DPO-PRO, a robust fine-tuning algorithm based on Direct Preference Optimization (DPO), which accounts for uncertainty in the preference distribution using a lightweight Distributionally Robust Optimization (DRO) formulation.
-Unlike prior DRO-based DPO methods, DPO-PRO is significantly less conservative.
+Unlike prior DRO-based variants, DPO-PRO focuses solely on uncertainty in preferences, avoiding unnecessary conservatism and incurring negligible computational overhead.
 We evaluate DPO-PRO on a real-world maternal mobile health program operated by the non-profit organization ARMMAN, as well as on standard alignment benchmarks.
@@
 Limited resource allocation is a core challenge in public health, where decision-makers must prioritize individuals for care or intervention under tight financial and operational constraints.
-In maternal healthcare, for instance, preventative care awareness programs have been shown to significantly reduce maternal mortality (Helpmum 2025; Jahan et al.
-2018; ARMMAN 2025), yet are often implemented by non-profit organizations that must serve large populations with limited resources.
+In maternal healthcare, for instance, preventative care awareness programs have been shown to significantly reduce maternal mortality (Helpmum; Jahan2018MomConnect; ARMMAN), yet are often implemented by non-profit organizations that must serve large populations with limited resources.
 To support such efforts, it is crucial to develop allocation policies that make effective sequential decisions based on individual-level attributes such as age, income, or health status.
 Reinforcement learning (RL) offers a powerful framework for learning allocation policies in complex, dynamic environments.
-A widely used model in this context is the Restless Multi-Armed Bandit (RMAB) (Niño-Mora 2023), a class of Markov Decision Processes (MDPs) well-suited for sequential decision-making under resource constraints.
-However, in RL, it is well-recognized that designing an appropriate reward function is critical (Singh, Lewis, and Barto 2009; Booth et al.
-2023).
+A widely used model in this context is the Restless Multi-Armed Bandit (RMAB) (math11071639), a class of Markov Decision Processes (MDPs) well-suited for sequential decision-making under resource constraints.
+However, in RL, it is well-recognized that designing an appropriate reward function is critical (singh09; Booth_Knox_Shah_Niekum_Stone_Allievi_2023).
 In public health applications, decision-makers’ priorities (e.g., which populations to prioritize) must be encoded into the reward function, so that resulting policy outcomes align with long-term public health goals.
@@
 As public health goals shift over time, this design process must be repeated, limiting the scalability and adaptability of RL-based solutions in dynamic, real-world environments.
-Figure 1: Overview of the implementation of DPO-PRO to public health applications such as the maternal health ARMMAN program in Section 5.2.
-Recently, several works explored using large language models (LLMs) to automate reward function design for sequential decision-making problems (Mirchandani, Karamcheti, and Sadigh 2021; Goyal, Niekum, and Mooney 2019; Carta et al.
-2022; Ma et al.
-2024; Behari et al.
-2024; Verma et al.
-2025).
+Recently, several works explored using large language models (LLMs) to automate reward function design for sequential decision-making problems (mirchan21; ijcai2019p331; carta2022eager; ma2024eureka; behari2024a; verma2025balancingactprioritizationstrategies).
 In these works, an LLM interprets a user’s high-level objective (e.g., desired policy outcomes) in natural language and outputs a reward function that guides RL policy training.
 This enables efficient workflow for developing human-aligned RL policies.
-Recent studies demonstrated that LLMs can effectively generate reward functions for RMAB problems in public health domains (Behari et al.
-2024; Verma et al.
-2025).
-Existing approaches often rely on inference-time optimization via self-reflection (Shinn et al.
-2023), where the LLM iteratively refines its outputs based on simulated feedback.
-While this strategy offers flexibility, it incurs substantial inference-time costs due to repeated rounds of simulation and reasoning.
-Furthermore, in non-profit settings, data sensitivity constraints often limit the use of large proprietary models, leading organizations to adopt smaller open-source alternatives that may be less capable off-the-shelf but can be finetuned.
-Because self-reflection is highly dependent on model strength, its effectiveness may be limited in these contexts.
-Fine-tuning offers a more practical alternative by directly improving model capabilities while avoiding the high inference costs and computational overhead that become prohibitive in large-scale public health deployments (Wu et al.
-2024), making it well-suited for resource-constrained domains such as public health (Zhao et al.
-2025).
+Recent studies demonstrated that LLMs can effectively generate reward functions for RMAB problems in public health domains (behari2024a; verma2025balancingactprioritizationstrategies).
+Existing approaches often rely on inference-time optimization via self-reflection (shinn2023reflexion), where the LLM iteratively refines its outputs based on simulated feedback.
+Although this strategy provides flexibility, it incurs substantial inference-time costs due to repeated rounds of simulation and reasoning.
+In many non-profit settings, data sensitivity constraints further limit the use of large proprietary models, pushing organizations toward open-source alternatives, which are often smaller and less capable.
+Because self-reflection is highly dependent on model strength, its effectiveness can be diminished in these environments.
+As a result, fine-tuning open-source models offers a more practical path: it directly improves their capabilities while avoiding the high inference costs and privacy issues in large-scale public health deployments (wu2024inference).
+This makes fine-tuning particularly suitable for resource-constrained domains such as public health (zhao2025foundation).
 Thus, we aim to move beyond inference-time methods and instead fine-tune an LLM to more effectively translate natural language preferences into high-performing reward functions for RMABs.
-A promising approach is Direct Preference Optimization (DPO) (Rafailov et al.
-2023), which enables LLM fine-tuning from pairwise preference annotations.
-Recent work has also leveraged LLM-as-a-judge to scale annotation as a proxy for human feedback (Zhang et al.
-2025; Tan et al.
-2024; Cui et al.
-2024; Zhu et al.
-2024).
+A promising approach is Direct Preference Optimization (DPO) (rafailov2023DPO), which enables LLM fine-tuning from pairwise preference annotations.
+Recent work has also leveraged LLM-as-a-judge to scale annotation as a proxy for human feedback (zhang2025aaai; tan-etal-2024-large; ultra2024; zhu2024starlingb).
 However, applying these techniques to the public health domain introduces significant challenges, particularly due to the presence of noise in the preference signal.
@@
 Second, reward design in RMABs demands complex reasoning, as the quality of a proposed reward function is not evaluated in isolation, but rather by the policy it induces, increasing the complexity of assessing preference.
-Third, real-world public health datasets are typically smaller, increasing the risk of misalignment due to overfitting and reward hacking, which is a common problem with reinforcement learning for human feedback methods such as DPO (Gheshlaghi Azar et al.
-2024; Amini, Vieira, and Cotterell 2024; Yang et al.
-2024; Xiao et al.
-2025).
+Third, real-world public health datasets are typically smaller, increasing the risk of misalignment due to overfitting and reward hacking, which is a common problem with reinforcement learning for human feedback methods such as DPO (pmlr-v238-gheshlaghi-azar24a; amini-etal-2024-direct; yang2024orthogonalfinetuningdirectpreference; xiao2025comprehensivesurveydirectpreference).
 However, public health is a high-stakes domain, where poorly aligned policies can lead to serious real-world consequences.
-Figure 2: DPO-PRO achieves comparable reward function performance to DLM on various ARMMAN maternal health tasks (Section 5.3) while providing (i) constant inference time scaling with only a finetuned LLama 8B LLM and (ii) removing external high-end LLM APIs for deployment.
-To tackle these challenges, we introduce DPO-PRO, a lightweight and distributionally robust enhancement of DPO (Rafailov et al.
-2023).
-Its application to public health settings is summarized in Figure 1.
-DPO-PRO incorporates uncertainty in the preference distribution using an efficient yet effective Distributionally Robust Optimization (DRO) formulation (Rahimian and Mehrotra 2022), ensuring that the fine-tuned policy remains reliable under imperfect preference annotation.
-Unlike prior DRO-based DPO methods (Wu et al.
-2025; Xu et al.
-2025; Mandal, Sasnauskas, and Radanovic 2025), our approach avoids excessive conservatism and only adds negligible computational overhead.
+To tackle these challenges, we introduce DPO-PRO, a lightweight and distributionally robust enhancement of DPO.
+DPO-PRO incorporates uncertainty in the preference distribution using an efficient yet effective Distributionally Robust Optimization (DRO) formulation (Rahimian_2022), ensuring that the fine-tuned policy remains reliable under imperfect preference annotation.
+Unlike prior DRO-based DPO methods (wu2025drdpo; xu2025distributionallyrobustdirectpreference; mandal2025distributionallyrobustreinforcementlearning), our approach avoids excessive conservatism and only adds negligible computational overhead.
 We further show that the resulting robust loss can be interpreted as a regularized DPO loss that penalizes model overconfidence and weak preference signals.
-We evaluate DPO-PRO in a simulated environment based on real-world maternal health application involving ARMMAN, a nonprofit organization that delivers preventative care messages via mobile health programs (ARMMAN 2025).
-Health workers can place a limited number of live service calls each week to boost beneficiary engagement.
-Prior work has modeled this problem using RMABs, where each beneficiary’s engagement evolves over time and the objective is to allocate $K$ live calls per week to maximize long-term engagement.
-Our experiments demonstrate that DPO-PRO is more robust to noisy preference supervision than both vanilla DPO and prior DRO-based approaches across a range of public health objectives.
-We also show that our method shows comparable performance with the Decision Language Model (DLM), a previous method based on self-reflection that was applied in the context of ARMMAN (Behari et al.
-2024), but requires significantly lower inference-time cost as shown in Figure 2.
-Additionally, our finetuned LLM has the benefit of removing the need for external LLM APIs during deployment, which is required when working with sensitive or private data.
+We evaluate DPO-PRO on the UltraFeedback benchmark (ultra2024) and on a simulated maternal health task based on ARMMAN’s real-world mobile health program (ARMMAN).
+In this setting, health workers can place only a limited number of live service calls each week to boost beneficiary engagement.
+Prior work models this problem using RMABs, where each beneficiary’s engagement evolves over time and the objective is to allocate a limited number of calls to maximize long-term engagement.
+Our experiments show that DPO-PRO is more robust to noisy preference annotation than vanilla DPO and prior DRO-based approaches, both on UltraFeedback and across a range of public health objectives.
+We further show that DPO-PRO achieves performance comparable to the Decision Language Model (DLM), a self-reflection–based method used in the ARMMAN setting (behari2024a), while requiring significantly lower inference-time cost (See Figure 1 in Section 5.2).
 2 Related Works
 A growing body of work has explored robustness in RLHF and DPO, particularly by leveraging DRO.
-Wu et al.
-(2025) introduce a DRO-based DPO formulation using a KL-divergence-based ambiguity set.
-Xu et al.
-(2025) extend this approach to both KL and Wasserstein distances, while Mandal, Sasnauskas, and Radanovic (2025) explore robustness under total variation distance, applying it to both DPO and standard two-stage RLHF.
+wu2025drdpo introduce a DRO-based DPO formulation using a KL-divergence-based ambiguity set.
+xu2025distributionallyrobustdirectpreference extend this approach to both KL and Wasserstein distances, while mandal2025distributionallyrobustreinforcementlearning explore robustness under total variation distance, applying it to both DPO and standard two-stage RLHF.
 In contrast to these DRO-based approaches, other works assume explicit corruption models.
-For example, Chowdhury, Kini, and Natarajan (2024) assume that preference labels are flipped according to a known noise rate, and Bukharin et al.
-(2024) model preferences using a Bradley–Terry (BT) framework (Bradley and Terry 1952) with added noise on the reward differences.
-Hong et al.
-(2024) use a DRO framework to develop adaptive loss functions for reward model learning, but focus is not on distributional robustness.
-Zhan et al.
-(2024) introduce robustness by constructing confidence sets over learned reward functions, followed by pessimistic policy optimization.
+For example, chowdhury2024 assume that preference labels are flipped according to a known noise rate, and bukharin2024robust model preferences using a Bradley–Terry (BT) framework (BT) with added noise on the reward differences.
+hong2024adaptive use a DRO framework to develop adaptive loss functions for reward model learning, but focus is not on distributional robustness.
+zhan2024provable introduce robustness by constructing confidence sets over learned reward functions, followed by pessimistic policy optimization.
 Broadly, existing robustness techniques fall into two categories: (1) methods relying on strong assumptions about noise models, which may be unrealistic in practice, and (2) DRO-based methods that avoid explicit corruption modeling but hedge against worst-case distributions.
 As discussed in Section 3.2, standard DRO approaches tend to be overly conservative and often require solving challenging min–max optimization problems.
-For tractability, prior methods rely on heuristic approximations (Wu et al.
-2025; Xu et al.
-2025), potentially undermining theoretical guarantees or practical robustness.
+For tractability, prior methods rely on heuristic approximations (wu2025drdpo; xu2025distributionallyrobustdirectpreference), potentially undermining theoretical guarantees or practical robustness.
 In contrast, we provide efficient approach DPO-PRO, with theoretical guarantees, and simultaneously illustrate its applicability in real world public health settings.
@@
 where $\sigma$ is the sigmoid function.
-Once the reward model is trained, a policy $\pi_{\theta}$ is optimized to maximize the expected reward while remaining close to a reference policy $\pi_{ref}$, typically via Proximal Policy Optimization (PPO) (Schulman et al.
-2017):
+Once the reward model is trained, a policy $\pi_{\theta}$ is optimized to maximize the expected reward while remaining close to a reference policy $\pi_{ref}$, typically via Proximal Policy Optimization (PPO) (schulman2017ppo):
 $\max_{\pi_{\theta}}\mathbb{E}_{(x,y)\sim\pi_{\theta}}\left[R_{\phi}(x,y)\right]-\beta\,\mathrm{KL}(\pi_{\theta}\,\|\,\pi_{\mathrm{ref}}),$
@@
 As discussed earlier, DPO is vulnerable to overfitting and noise in the data.
-To address this, recent work has proposed distributionally robust objectives of the form $\max_{P\in\mathcal{Q}(\mathcal{D})}\mathcal{L}_{{P}}(\pi_{\theta}),$ where $\mathcal{Q}(\mathcal{D})$ denotes an ambiguity set centered around the empirical distribution (data) $\mathcal{D}$ and the loss $\mathcal{L}_{{P}}$ is the DPO loss evaluated under the perturbed distribution $P$ rather than the original data distribution (Xu et al.
-2025; Wu et al.
-2025).
+To address this, recent work has proposed distributionally robust objectives of the form $\max_{P\in\mathcal{Q}(\mathcal{D})}\mathcal{L}_{{P}}(\pi_{\theta}),$ where $\mathcal{Q}(\mathcal{D})$ denotes an ambiguity set centered around the empirical distribution (data) $\mathcal{D}$ and the loss $\mathcal{L}_{{P}}$ is the DPO loss evaluated under the perturbed distribution $P$ rather than the original data distribution (xu2025distributionallyrobustdirectpreference; wu2025drdpo).
 Conservatism and Computational Costs
@@
 In our setting, the prompt $x$ represents a desired policy outcome specified by a human prompter, for example, Focus on the young mothers by age and also focus on those with low income.
-See Figure 1.
-The responses $y_{1}$ and $y_{2}$ correspond to candidate reward functions for RMABs (details of the RMAB formulation are provided in the Appendix).
+The responses $y_{1}$ and $y_{2}$ correspond to candidate reward functions for RMABs.
 The role of the LLM-annotator is to judge which of the two reward functions, $y_{1}$ or $y_{2}$, better aligns with the objective expressed in the prompt.
 A key challenge in this task is that the quality of a reward function cannot be assessed in isolation.
-Rather, it must be evaluated based on the policy it induces when optimized, i.e., how effectively the resulting policy fulfills the prompter’s (often ambiguous) intent.
+Rather, it must be evaluated based on the policy it induces when optimized, i.e., how effectively the resulting policy fulfills the user’s (often ambiguous) intent.
 This requires the annotator to infer the downstream effects of each reward function.
@@
 In this section, we introduce DPO-PRO, a distributionally robust version of DPO that specifically targets distributional shifts or other forms of noise in the preference distribution.
+The complete proofs for all theorems are provided in the extended version of this paper (kim2025dpopro), which includes the full technical appendix.
 4.1 Uncertainties in the Preference Distribution
@@
 As more data is collected, the uncertainty in this part of the data is expected to diminish.
-Noise in the prompts and responses becomes even less of an issue under iterative DPO methods, which are increasingly adopted in practice (Xiong et al.
-2024; Cen et al.
-2025; Xie et al.
-2025).
+Noise in the prompts and responses becomes even less of an issue under iterative DPO methods, which are increasingly adopted in practice (xiong2024iterative; cen2025valueincentivized; xie2025exploratory).
 In contrast, we argue that the preference distribution $c\sim p^{*}$ is inherently noisy in practice, and collecting more data is unlikely to resolve the underlying uncertainty in human preference.
@@
 Our formulation avoids unnecessary robustification over the joint distribution on $(x,y_{1},y_{2})$, and assumes that this part of the data is reliable.
-Furthermore, unlike standard DRO methods that can add significant computational burden (Rahimian and Mehrotra 2022), our formulation introduces negligible additional cost as shown next.
+Furthermore, unlike standard DRO methods that can add significant computational burden (Rahimian_2022), our formulation introduces negligible additional cost as shown next.
 4.2 Computing Worst-case Distribution
@@
 Note that we do not formally differentiate through the inner maximization in Eq (1) with respect to $\theta$.
-However, Danskin’s theorem (Bertsekas 1999) justifies our approach.
+However, Danskin’s theorem (Bertsekas99Np) justifies our approach.
 For a fixed $\theta$, we may solve the inner maximization and directly substitute the resulting worst-case $\hat{p}$ into the gradient expression.
@@
 The DRO adjustment described above requires a soft score $q(y_{1}\succ y_{2}|x)$, rather than a binary annotation for each prompt and response pair $(x,y_{1},y_{2})$.
-In this work we focus on settings where preferences are evaluated by LLM judges.
-Soft scores can be obtained from LLM judges in several ways: (1) producing multiple binary preference judgments through repeated queries and averaging the outcomes; (2) directly outputting a numerical preference score via appropriate prompting; or (3) extracting the log-probabilities assigned to tokens representing the preferences (e.g., 1 for $y_{1}\succ y_{2}$ and -1 for $y_{2}\succ y_{1}$), and applying a softmax transformation to estimate a preference distribution (Lee et al.
-2024).
-We provide further discussion on annotation formats and the connection to robustness in the Appendix.
+Soft scores can be obtained from LLM judges in several ways: (1) producing multiple binary preference judgments through repeated queries and averaging the outcomes; (2) directly outputting a numerical preference score via appropriate prompting; or (3) extracting the log-probabilities assigned to tokens representing the preferences (e.g., 1 for $y_{1}\succ y_{2}$ and -1 for $y_{2}\succ y_{1}$), and applying a softmax transformation to estimate a preference distribution (lee2024).
+When using a off-the-shelf reward model, the scalar outputs of the reward model can be converted into a pairwise preference probability using the BT model.
+We provide further discussion on annotation formats and the connection to robustness in (kim2025dpopro).
 5 Experiments
@@
 First, we validate the methodology on a standardized benchmark for preference-based LLM finetuning.
-Next, we adapt this experiment for the ARMMAN maternal health dataset (ARMMAN 2025) in the following section.
+Next, we adapt this experiment for the ARMMAN maternal health dataset (ARMMAN) in the following section.
+The complete experimental details are provided in the extended version of this paper (kim2025dpopro).
 5.1 Benchmarks on General Alignment Datasets
-We use the recent UltraFeedback alignment benchmark dataset (Cui et al.
-2024), consisting of 60,000 top-quality human pairwise preference annotations that cover a wide variety of tasks and instruction types, rendering it a broad benchmark for examining alignment research.
+We use the recent UltraFeedback alignment benchmark dataset (ultra2024), consisting of 60,000 top-quality human pairwise preference annotations that cover a wide variety of tasks and instruction types, rendering it a broad benchmark for examining alignment research.
 Noise in the training data
@@
 We use $q^{*}$ denote the true preference distribution used at test time.
-For our experiments, we define $q^{*}$ using the state-of-the-art preference reward model Eurus 7B(Yuan et al.
-2025), which has been trained for the UltraFeedback dataset (Cui et al.
-2024).
+For our experiments, we define $q^{*}$ using the state-of-the-art preference reward model Eurus 7B(yuan2025eurus), which has been trained for the UltraFeedback dataset (ultra2024).
 We compute $q^{*}$ using the BT model.
@@
 During training, we have access to $q_{\alpha}$ but not to $q^{*}$, with $\alpha$ being unknown.
-We include additional experiments with alternative forms of noise in Appendix.
 Training details
-We base our experiments on the recent Phi 3-mini 3B model without instruction (Abdin et al.
-2024).
+We base our experiments on the recent Phi 3-mini 3B model without instruction (abdin2024phi3).
 We initially train for two epochs of supervised fine tuning (SFT), followed by one epoch of DPO, with the baselines discussed below.
 All experiments are performed on a Nvidia A100 GPU 40 GB with batch size 2.
-Refer to the technical appendix for full details on the training procedure such as weight decay, learning rate scheduler, low-rank adaptation parameters, quantization, etc and experiments conducted on the Llama 3 8B model (Grattafiori et al.
-2024), not shown here due to space limitations.
 Baselines
-We compare DPO-PRO against vanilla DPO (Rafailov et al.
-2023) and the two robust baselines: rDPO (Chowdhury, Kini, and Natarajan 2024) and DrDPO (Wu et al.
-2025).
-We implement DPO-PRO with a value of $\rho=0.1$ for the $\chi^{2}$-divergence.
-We train for two noise levels ($\alpha=0.1$ and $\alpha=0.3$), termed low and high respectively, and also for the noiseless case $\alpha=0$.
-Evaluation metrics
-After the DPO step, we generate a response for every prompt $x$ in the evaluation dataset $y_{g}\sim\pi_{\theta}(x)$.
-Then, we compute the reward $R_{\phi}(x,y_{g})$.
-We focus on two evaluation metrics, the win rate, which is computed by comparing the generated prompt with the chosen response $y_{c}$ in the dataset as $\texttt{win\_rate}=\frac{1}{N_{\mathrm{eval}}}\sum_{i=1}^{N_{\mathrm{eval}}}\mathbb{I}(R_{\phi}(x,y_{g})>R_{\phi}(x,y_{c})),$ where $\mathbb{I}$ is the indicator function.
-The next metric is the generation evaluation reward, computed simply as $\texttt{eval\_reward}=\frac{1}{N_{\mathrm{eval}}}\sum_{i=1}^{N_{\mathrm{eval}}}R_{\phi}(x,y_{g})$.
+We compare DPO-PRO against the standard DPO (rafailov2023DPO) and the prior distributionally robust variant DrDPO (wu2025drdpo) using their best hyperparameters suggested by the authors.
+We evaluate DPO-PRO under several values of the robustness parameter $\rho$ based on the $\chi^{2}$-divergence.
+Models are trained under two levels of label-flip noise, $\alpha=0.3$ (low) and $\alpha=0.6$ (high), as well as the noiseless setting $\alpha=0$.
+Evaluation
+After the fine-tuning step, we generate a response $y_{g}\sim\pi_{\theta}(x)$ for every prompt $x$ in the evaluation dataset and compute its reward $R_{\phi}(x,y_{g})$.
+We consider two evaluation metrics.
+The first is the win rate, computed by comparing the generated response with the chosen response $y_{c}$ in the dataset using an LLM-as-a-judge approach with GPT-4o-mini (openai_gpt4omini_2024).
+To mitigate position bias, we randomly flip the order of the two responses presented to the judge.
+The second metric, evaluation reward, is computed as the average reward score defined as $\frac{1}{N_{\mathrm{eval}}}\sum_{i=1}^{N_{\mathrm{eval}}}R_{\phi}(x,y_{g}).$
 Results
-Table 1(b) presents the results.
-Across all noise levels, DPO-PRO consistently achieves highest performance in both win rate and evaluation reward.
-Notably, it outperforms all baselines even in noise-free setting, highlighting robustness without sacrificing standard performance.
-We found it surprising that performance of all methods increased with additional noise.
-This can be explained by the fact that DPO in fixed (off-policy) data is known to be highly sensitive to reward hacking (Gheshlaghi Azar et al.
-2024; Amini, Vieira, and Cotterell 2024; Yang et al.
-2024; Xiao et al.
-2025).
-An analogous result for DPO-based methods can be found in Wu et al.
-(2025); Bukharin et al.
-(2024).
-The noise seemingly acts as a regularization, preventing model from overfitting to spurious patterns in training data and instead learning more generalizable features.
-This phenomenon is consistent with recent theoretical work showing controlled noise injection can improve generalization by encouraging simpler decision boundaries (Li et al.
-2025).
-Table 1: Results on UltraFeedback under label-flip noise. Experiments run with Phi-3-mini. Mean and standard error over 2000 evaluation samples are shown.
-(a) Win rate ($\uparrow$)
-Method No noise Low noise (0.1) High noise (0.3)
-[…]
-[…]
-[…]
-[…]
-(b) Generation Evaluation Reward ($\uparrow$)
-Method No noise Low noise (0.1) High noise (0.3)
-[…]
-[…]
-[…]
-[…]
+Table 1 summarizes the results.
+For the LLM-as-a-judge evaluation, the different DPO-PRO variants (each corresponding to different $\rho$ value) tend to underperform the baselines in the low-noise setting, but their relative performance improves as noise increases.
+Under high noise, all DPO-PRO variants outperform both vanilla DPO and DrDPO.
+The effect of $\rho$ follows the expected pattern: smaller $\rho$ values perform well when noise is low but degrade more quickly as noise increases, while larger $\rho$ values exhibit more stability.
+Overall, the DPO-PRO family shows improved robustness as the preference noise becomes larger.
+For metrics based on the reward model, DPO-PRO outperforms the baselines even in the noiseless setting, except for the case with $\rho=0.1$.
+As expected, performance declines for all methods as noise increases.
+In the low-noise regime, however, the DPO-PRO variants perform slightly worse than the baseline, but they regain and ultimately surpass baseline performance as the noise becomes even larger.
+Note that this does not contradict the DRO objective, because the true optimization target for fine-tuning is reward plus KL regularization, while our evaluation here relies solely on reward-model scores.
+Moreover, these scores are only a proxy for quality, particularly for out-of-distribution responses, and are limited by the reward model’s own generalization capabilities.
+(a) Win rate (%) based on LLM-as-a-judge evaluation ($\uparrow$).
+Method No noise Low noise High noise
+DPO-PRO ($\rho=0.008$) 43.24 35.17 28.75
+DPO-PRO ($\rho=0.03$) 41.71 33.33 28.98
+DPO-PRO ($\rho=0.1$) 40.33 33.54 27.97
+DPO 43.77 36.19 26.97
+DrDPO 44.85 38.84 25.51
+(b) Evaluation reward ($\uparrow$).
+Method No noise Low noise High noise
+DPO-PRO ($\rho=0.008$) 923.70 713.00 446.65
+DPO-PRO ($\rho=0.03$) 912.65 695.00 499.59
+DPO-PRO ($\rho=0.1$) 781.28 656.32 499.36
+DPO 877.01 729.12 400.44
+DrDPO 897.67 753.77 419.05
+Table 1: Results on UltraFeedback under label-flip noise.
 5.2 Benchmarks on Maternal Health Data
-We apply our proposed method in the context of mobile health program run by ARMMAN (2025).
-Specifically, we have anonymized beneficiary listenership data from a quality-improvement study conducted by ARMMAN [^1] The authors have been access to this restricted dataset under a data usage agreement..
+We apply our proposed method in the context of mobile health program run by ARMMAN.
+We use real-world anonymized beneficiary listenership data from a quality-improvement study conducted by ARMMAN [^1] The authors have been access to this restricted dataset under a data usage agreement..
 The data consists of registration information of beneficiaries (mothers) containing their socio-demographic features of age, income education as well as their preferences on call slot times and language.
@@
 We model the problem of allocating live service calls as a RMAB problem.
-For all our experiments, we consider a sample of population with $2100$ beneficiaries and $K=210$ as the budget, representing the number of live service calls that can be made every week.
-Exact policy optimization for RMABs is generally known to be intractable (Papadimitriou and Tsitsiklis 1987).
-However, Whittle index policy (Whittle 1988) is a popular heuristic to obtain an asymptotically optimal policy.
-Please refer also to the technical appendix for additional details.
+For all our experiments, we consider a population with $2100$ beneficiaries (i.e., 2100 arms) and the number of live service calls that can be made every week is limited to 210.
+We use the Phi 3-Mini 3B model, and the noise-injection and training procedures follow the same setup as in Section 5.1.
+Here, we experiment only with the noiseless and low-noise settings to better reflect realistic conditions, rather than injecting artificially large noise.
 Preference Dataset Construction
-We generate a preference dataset of health worker prioritization commands, preferred, and rejected reward functions as shown in Figure 1.
-Since it is costly to generate this dataset through human annotation, we use an LLM judge with ChatGPT 4o-mini (OpenAI 2024) as follows: i) query LLM-judge to obtain 20 candidate reward functions that align with the prioritization command, ii) for each candidate reward function, solve RMAB problem using Whittle index method and generate trajectory outputs, iii) sample 50 pairs of reward functions from this set and query the LLM-judge to select preferred reward function, iv) perform this query 10 times to estimate the LLM’s uncertainty over preferences.
-Finally, we obtain a dataset of 9500 preferred and rejected reward function responses over 190 prioritization commands (see Appendix for more details).
-Noise injection, training details, and baselines
-We adopt the same noise injection scheme to obtain $q_{\alpha}$ as described in Section 5.1 and Eq (4).
-We fine-tune a Llama 3 (8B) model, keeping the learning rate, DPO reward temperature, $\chi^{2}$-divergence radius, and other hyperparameters identical to those used in the UltraFeedback benchmark experiments.
-Given poor performance of rDPO in Table 1(b), we omit this baseline in this experiment.
+We generate a preference dataset of health worker prioritization commands, preferred, and rejected reward functions.
+Since it is costly to generate this dataset through human annotation, we use an LLM judge with ChatGPT 4o-mini (openai_gpt4omini_2024) as follows: i) query LLM-judge to obtain 20 candidate reward functions that align with the prioritization command, ii) for each candidate reward function, solve RMAB problem using Whittle index method (whittle1988restless) and generate trajectory outputs (see (math11071639) or the Appendix in (kim2025dpopro) for the explanation on the Whittle index), iii) sample 50 pairs of reward functions from this set and query the LLM-judge to select preferred reward function, iv) perform this query 10 times to estimate the LLM’s uncertainty over preferences.
+Finally, we obtain a dataset of 9500 preferred and rejected reward function responses over 190 prioritization prompts.
+Baselines
+In addition to the DPO variants, we also compare against the Decision Language Model (DLM), a self-reflection–based inference-time approach introduced in (behari2024a) using Gemini 2.0.
+Note that DLM is not fine-tuned on the training data, so this method has no notion of noise.
+For each prompt, we simply apply the out-of-the-box Gemini model and run the self-reflection procedure described in (behari2024a).
+We use Gemini 2.0, as in the original work, rather than the open-source Phi-3B model used for fine-tuning.
+This is to assess how our method compares against a state-of-the-art closed model that a typical user would rely on, further strengthened by self-reflection.
 Evaluation
-To evaluate the quality of a generated reward function, we follow a procedure analogous to the dataset creation: i) solve the resulting RMAB problem using the Whittle Index method to generate policy trajectories, ii) use LLM judge (GPT-4o-mini) to evaluate these policy outcomes.
-Specifically, we compute the win rate by comparing the policy induced by our generated reward function to the policy obtained from the reward function in the training data.
-Table 2: Examples of reward functions produced by Vanilla-DPO and DPO-PRO when trained under noise. DPO-PRO produces more conservative preference interpretations.
+To evaluate the quality of each generated reward function, we follow a procedure parallel to the dataset construction.
+For every prompt in the evaluation set, we generate a reward function using a given fine-tuning method and obtain the corresponding RMAB policy using the Whittle Index.
+We repeat the same procedure using the chosen reward function associated with that prompt in the evaluation data, producing a reference policy for comparison.
+We compare these policies using two complementary methods.
+First, because this task provides access to ground-truth human-designed reward functions from behari2024a, we can directly evaluate RMAB policy quality by computing the achieved reward under this ground-truth function.
+We compare the average ground-truth reward obtained by the policy induced by the fine-tuned LLM with that obtained by the policy from the evaluation data.
+Second, we also perform an LLM-as-a-judge evaluation using GPT-4o-mini to compute win-rate as in Section 5.1.
+The judge compares the resulting trajectories of the RMAB policies, i.e., the resource allocation outcomes across the demographic data, to determine which policy is preferred.
+Results
+We report the results in Table 2.
+For the ground-truth reward metric, all variants of DPO-PRO outperform both DrDPO and Vanilla DPO in both the noiseless and high-noise settings.
+Their performance is slightly lower than DLM, but this gap must be interpreted in context: DLM requires (1) access to a high-end proprietary LLM API at deployment time and (2) substantially higher inference-time cost, as we show below.
+For the win-rate metric, the results are more mixed.
+The DPO-PRO variants generally outperform DrDPO and Vanilla DPO in the low-noise setting, with the exception of the $\rho=0.03$ variant.
+In particular, DPO-PRO ($\rho=0.1$) outperforms both DrDPO and Vanilla DPO in both the noiseless and low-noise settings, while DPO-PRO ($\rho=0.3$) performs similarly to these baselines under no noise and clearly exceeds them under low noise.
+As with the ground-truth reward metric, all DPO-PRO variants perform below DLM.
+One interesting observation is that, unlike the clean trends in Section 5.1, the performance across different $\rho$ values and noise levels in the ARMMAN task does not yield a simple or monotonic pattern.
+For example, larger $\rho$ values are not consistently more conservative, and performance even improves as noise is injected.
+Although pinpointing the exact cause is difficult, several factors likely contribute to this behavior.
+First, unlike the curated benchmark data in Section 5.1, the ARMMAN dataset is substantially noisier.
+Preference labels generated by LLM judges are difficult to obtain reliably for this task (see Section 3), reducing the quality of the signal available during fine-tuning.
+In other words, the noiseless setting here does not necessarily correspond to a clean preference signal, since the underlying annotations are already highly noisy.
+Second, DPO and related off-policy preference-based methods are known to be sensitive to overfitting and reward hacking when trained on fixed datasets (pmlr-v238-gheshlaghi-azar24a; amini-etal-2024-direct; yang2024orthogonalfinetuningdirectpreference; xiao2025comprehensivesurveydirectpreference).
+In such settings, injected noise can act as a regularizer: it discourages the model from fitting spurious or inconsistent preference patterns and promotes more stable decision rules.
+Similar trends, where performance improves under moderate noise, have been reported in (wu2025drdpo; bukharin2024robust).
+This interpretation is also consistent with recent theoretical work showing that controlled noise injection can improve generalization by implicitly encouraging simpler and more robust decision boundaries (li2025towards).
+Inference Time Comparison
+Figure 1 compares the inference time of DPO-PRO (with $\rho=0.1$) and DLM.
+Unlike DLM, whose inference time scales linearly with the number of arms due to repeated simulation of resource-allocation policy and self-reflection, DPO-PRO incurs a constant inference cost regardless of problem size.
+This efficiency gap becomes increasingly significant in large-scale settings, making DPO-PRO more scalable.
+While overall performance is comparable or mixed across evaluation metrics, minimal inference-time overhead makes DPO-PRO especially well-suited for real-world deployment.
+Examples
+To further illustrate the differences between vanilla DPO and DPO-PRO (with $\rho=0.1)$, we present response outputs from both models on two representative tasks from the evaluation dataset (Table 3).
+Many tasks contain attributes that can be interpreted in multiple ways.
+For instance, the concept of “young beneficiaries” can be identified directly through the age feature, but also indirectly via lower education level or income.
+Similarly, a task that expresses a preference for “midday calls” could refer explicitly to the 12:30–3:00 PM time slot, but might also include the 10:30–12:30 PM window, depending on interpretation.
+In such cases, we observe that vanilla-DPO favors uncommon interpretations of ambiguous attributes based on spurious correlations in noisy preference data.
+In contrast, DPO-PRO adopts a cautious strategy by selecting narrower definition, aiming to be robust to uncertainty in the preference signal at evaluation.
+(a) Average Ground Truth Reward ($\uparrow$).
+Method No noise Low noise
+DLM $6073.04$ NA
+DPO $5465.36$ $5707.09$
+DrDPO $5480.06$ $5675.17$
+DPO-PRO ($\rho=0.008$) $5530.58$ $5832.66$
+DPO-PRO ($\rho=0.03$) $5557.9$ $5790.8$
+DPO-PRO ($\rho=0.1$) $5677.76$ $5807.83$
+DPO-PRO ($\rho=0.3$) $5797.1$ $5904.7$
+(b) Win rate (%, $\uparrow$).
+Method No noise Low noise
+DLM 24.82 NA
+DPO 15.38 13.84
+DrDPO 15.89 12.13
+DPO-PRO ($\rho=0.008$) 14.18 14.87
+DPO-PRO ($\rho=0.03$) 13.5 13.5
+DPO-PRO ($\rho=0.1$) 17.7 14.35
+DPO-PRO ($\rho=0.3$) 15.7 17.8
+Table 2: Results on ARMMAN environment under no-noise and low-noise settings.
+Figure 1: Inference-time comparison of DPO-PRO and DLM across different population sizes (i.e., number of arms) in the underlying RMAB.
 Task 1: Prefer both young and elderly beneficiaries
@@
 DPO-PRO s + 3 * (12_30-3pm and NGO_registered)
-Figure 3: Winrates for RMAB environment under label flipping noise. We present more results in the Appendix.
-Results
-Table 3 shows the results.
-In zero-noise setting, vanilla-DPO outperforms robust methods.
-However, as noise level increases, DPO-PRO outperforms vanilla DPO and DrDPO.
-We also perform a task-level analysis of winrates in Figure 3.
-We note that across all tasks, under the presence of noise, robust DRO methods are better than vanilla-DPO.
-Especially, in 5 out of 8 evaluation task categories, DPO-PRO outperforms baselines under the presence of noise.
-This signifies the robustness of our approach.
-To further illustrate the differences between vanilla DPO and DPO-PRO, we present response outputs from both models on two representative tasks from the evaluation dataset (Table 2).
-Many tasks contain attributes that can be interpreted in multiple ways.
-For instance, the concept of “young beneficiaries” can be identified directly through the age feature, but also indirectly via lower education level or income.
-Similarly, a task that expresses a preference for “midday calls” could refer explicitly to the 12:30–3:00 PM time slot, but might also include the 10:30–12:30 PM window, depending on interpretation.
-In such cases, we observe that vanilla-DPO favors uncommon interpretations of ambiguous attributes based on spurious correlations in noisy preference data.
-In contrast, DPO-PRO adopts a cautious strategy by selecting narrower definition, aiming to be robust to uncertainty in the preference signal at evaluation.
-5.3 Comparison with Decision-Language Model
-Table 4 compares DPO-PRO with the Decision Language Model (DLM), a self-reflection-based approach introduced in (Behari et al.
-2024), using two evaluation metrics: (1) win rate, determined by an external LLM judge (GPT-4o-mini), and (2) reward, computed using ground-truth human-designed reward functions.
-While DLM slightly outperforms DPO-PRO on both metrics, the results indicate that DPO-PRO achieves comparable alignment quality despite its simplicity and significantly lower inference-time cost.
-Figure 2 provides detailed comparison of inference-time cost between DPO-PRO and DLM.
-Unlike DLM, whose inference cost scales linearly with number of arms due to repeated simulation and self-reflection, DPO-PRO incurs a constant inference cost regardless of problem size.
-This efficiency gap becomes increasingly significant in large-scale settings, making DPO-PRO more scalable.
-While overall performance is comparable or mixed across evaluation metrics, minimal inference-time overhead makes DPO-PRO especially well-suited for real-world deployment.
-Table 3: Winrate (%, $\uparrow$) comparision for ARMMAN environment for different methods under no noise and high noise settings, aggregated across all tasks.
-Method Vanilla-DPO DrDPO DPO-PRO
-No Noise $\bf 55\pm 2.3$ $40\pm 1.85$ $35.3\pm 1.83$
-High Noise (0.3) $45\pm 2.03$ $54.8\pm 1.84$ $\bf 63.9\pm 1.94$
-Table 4: Comparison of Winrate and Reward
-Method Winrate Ground Truth Reward
-DLM $35.5\pm 5.02$ $6073.04\pm 88.5$
-DPO-PRO $32.2\pm 4.92$ $5687.3\pm 112.3$
+Table 3: Examples of reward functions produced by Vanilla-DPO and DPO-PRO when trained under noise. DPO-PRO produces more conservative preference interpretations.
 6 Conclusion
-Motivated by reward function design tasks in RMABs for public health, we introduced DPO-PRO, a robust fine-tuning algorithm that enhances DPO by explicitly modeling uncertainty in preference distributions through a lightweight DRO formulation.
-Our method is particularly well-suited for high-stakes decision-making settings like public health, where preference annotations are inherently noisy and often limited.
-Compared to existing DRO-based approaches, DPO-PRO avoids excessive conservatism while providing strong theoretical justification as a regularized variant of DPO.
-Through extensive evaluation on a real-world maternal health program and standard alignment benchmarks, we demonstrated that DPO-PRO improves robustness, and achieves competitive performance to inference-time methods such as self-reflection at a fraction of the computational cost.
-Additionally, contrary to inference-time methods, DPO-PRO removes the dependency on external LLM APIs at deployment which is essential when working with sensitive or private data.
+Motivated by reward-function design in RMABs for public health, we introduced DPO-PRO, a robust fine-tuning algorithm that extends DPO by explicitly modeling uncertainty in preference distributions through a lightweight DRO formulation.
+Our method is particularly well-suited for high-stakes decision-making settings such as public health, where preference annotations are inherently noisy and often limited.
+Compared to existing DRO-based approaches, DPO-PRO avoids excessive conservatism while offering a theoretical interpretation as a regularized variant of DPO.
+Through extensive evaluation on real-world public health data and standard alignment benchmarks, we showed that DPO-PRO improves robustness and achieves performance competitive with inference-time methods at a fraction of their computational overhead.
+Moreover, compared with the previous inference-time approach, DPO-PRO eliminates the need for external LLM APIs at deployment, an essential property when handling sensitive or private data.
 These results highlight the potential of DPO-PRO as a scalable and reliable tool for learning aligned reward functions in real-world policy learning tasks for public health.
-Appendix A Codebase
-We release the codebase for all our experiments on finetuning LLMs here: https://anonymous.4open.science/r/DPO-Pro-A33D/README.md
-Appendix B Extensions and Generalizations
-B.1 Beyond the Chi-squared Divergence
-We use the chi-squared divergence to define the ambiguity set due to the simplicity of its resulting closed-form solution.
-However, alternative divergence measures can also be used and are likely to offer comparable computational efficiency as ours.
-This is because the per-sample optimization problem in Eq (2) remains a one-dimensional problem with linear objective.
-For example, using the KL divergence leads to a constraint of the form: $p\log\frac{p}{q}+(1-p)\log\frac{1-p}{1-q}\;\leq\;\rho$.
-In this case, the worst-case $\hat{p}$ can be computed by solving the equation $p\log\frac{p}{q}+(1-p)\log\frac{1-p}{1-q}\;=\;\rho$, which can be efficiently handled using a simple one-dimensional root-finding method.
-B.2 Obtaining the Soft-score $q$
-In many practical applications beyond ours, only a single binary annotation might be available per $(x,y_{1},y_{2})$.
-In this case, we effectively observe only one sample from the underlying Bernoulli distribution $q(y_{1}\succ y_{2}|x)$, which makes estimating $q$ challenging.
-However, our method naturally accommodates binary annotations, which can be especially useful when part of the dataset is labeled by humans.
-For instance, a healthcare worker involved in the annotation process may choose to directly express their preferences, in which case requesting a binary label is more appropriate.
-We can slightly relax the definition of chi-squared divergence by rewriting the constraint as $(p-q)^{2}\leq\rho\,q(1-q),$ which reduces to $\hat{p}=q$ when $q\in\{0,1\}$.
-As a result, when hard binary labels are used, the DRO loss reduces to the standard DPO loss, introducing no additional regularization.
-This effectively means that we regularize only on LLM-generated annotations, while treating human-provided labels as ground truth.
-To inject robustness in such cases, we can smooth the binary label by replacing the label $1$ with $1-\epsilon$, and 0 with $\epsilon$, where $\epsilon$ is a small hyperparameter.
-The DRO formulation can then be applied to this smoothed label.
-This makes the uncertainty-weighted coefficient uniform across all pairs (since $q$ only takes values $1-\epsilon$ or $\epsilon$), but the $(\ell_{1}-\ell_{-1})$ term still encourages the model to reduce overconfidence in its current preference.
-Our method also applies to the case where off-the-shelf reward models are used for DPO (Guo et al.
-2024; Ding et al.
-2024; Lee et al.
-2025; Gupta et al.
-2025).
-In this setting, acquiring soft score $q(y_{1}\succ y_{2}|x)$ is straightforward.
-The scalar outputs of the reward model can be converted into a pairwise preference probability using the BT model.
-However, these preference probabilities can be noisy or uncalibrated.
-We discuss alternative annotation formats and how our approach adapts robustness to each setting in the Appendix.
-B.3 Uncertainty-Weighted Coefficient
-We visualize this uncertainty-weighted coefficient for various $\rho$ values in Figure 4.
-Figure 4: Visualization of the uncertainty-weighted coefficient for various values of $\rho$.
-This coefficient attains its maximum near $q=0.5$ and decreases as $q\to 0$ or $q\to 1$.
-For small values of $\rho$, the maximum is achieved exactly at $q=0.5$, where $\sqrt{\rho\,q(1-q)}$ peaks.
-However, for larger $\rho$, the maximum occurs at the intersection point between the curves 1q1-q and $\sqrt{\rho\,q(1-q)}$.
-This is because $q+\sqrt{\rho\,q(1-q)}$ (the worst-case distribution under the chi-squared divergence constraint) may exceed 1 when $\rho$ is large, causing the worst-case distribution $\hat{p}$ to be clipped at 1.
-In such cases, the adversary’s perturbation $\hat{p}-q$ can be greater when $q<0.5$, leading the penalty term to peak at a value of $q$ smaller than 0.5.
-Appendix C Omitted Proofs
-C.1 Proof of Proposition 4.1
-The result follows from Danskin’s theorem (Bertsekas 1999).
-To apply the theorem, we must verify that the inner maximization problem in Eq.
-(2) has a compact feasible set and admits a unique optimal solution.
-To establish compactness, we need to show that the feasible set is both bounded and closed.
-Boundedness is immediate since $p\in[0,1]$.
-Closedness follows from the fact that the constraint function $p\mapsto(p-q)^{2}-\rho q(1-q)$ is continous in $p$.
-Hence, the set $(p-q)^{2}\leq\rho q(1-q)$ is closed.
-Since $p\in[0,1]$ is also closed and the intersection of closed sets is closed, the feasible set is closed.
-Uniqueness of the optimal solution $\hat{p}$ is also straightforward, since the objective is one-dimensional linear function in $p$.
-C.2 Proof of Proposition 4.2
-We begin by assuming the case $\ell_{1}\geq\ell_{-1}$. Then,
-$\mathcal{L}_{\text{DRO}}$ $=\hat{p}\,\ell_{1}+(1-\hat{p})\,\ell_{-1}$
-$=q\,\ell_{1}+(1-q)\,\ell_{-1}+(\hat{p}-q)(\ell_{1}-\ell_{-1})$
-$=\mathcal{L}_{\text{DPO}}+\underbrace{\min\left\{1-q,\;\sqrt{\rho\,q(1-q)}\right\}}_{\text{Uncertainty-Weighted Coefficient}}(\ell_{1}-\ell_{-1}).$
-For the opposite case, when $\ell_{-1}>\ell_{1}$, the derivation proceeds symmetrically.
-$\mathcal{L}_{\text{DRO}}$ $=\hat{p}\,\ell_{1}+(1-\hat{p})\,\ell_{-1}$
-$=\mathcal{L}_{\text{DPO}}+(\hat{p}-q)(\ell_{1}-\ell_{-1})$
-$=\mathcal{L}_{\text{DPO}}-\min\left\{q,\,\sqrt{\rho\,q(1-q)}\right\}(\ell_{1}-\ell_{-1})$
-$=\mathcal{L}_{\text{DPO}}+\min\left\{q,\,\sqrt{\rho\,q(1-q)}\right\}(\ell_{-1}-\ell_{1}).$
-Appendix D Multi-armed Restless Bandits
-A Multi-Armed Restless Bandit problem is characterized by N Markov Decision processes, each defined by the tuple $\{\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{T},\gamma\}$.
-Here, $\mathcal{S}$ represents the set of states, $\mathcal{A}$ represents the set of actions which are binary in $\{0,1\}$, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ represents the transition probability from current state, under an action to the next state.
-$\mathcal{R}:\mathcal{S}\to\mathbb{R}$ represents a reward function mapping state to a real number, and $\gamma$ represents the discounting factor.
-We also have a budget $K$ which represents the number of arms that can be chosen at every timestep.
-The objective of the Multi-Armed Restless Bandit problem then is to decide a vector of actions $\textbf{A}^{t}$ representing actions $a_{i}^{t}\in\mathcal{A}$ for every arm i in time step $t$ such that we maximize long term reward under the budget constraint:
-$J(s)=\max_{\textbf{A}\in\mathcal{A}}\left(\sum_{i=1}^{N}R_{i}(s_{i})+\gamma\mathbb{E}[J(s^{\prime})\mid s,\textbf{A}]\right)$
-$s.t.\sum_{i=1}^{i=N}A_{i}^{t}\leq K\;\forall\;t=1\;\text{to}\;T$
-where $s_{t}\sim T(s_{t-1},a_{t},.)$
-Appendix E Experimental Details
-E.1 Additional Comparison with Baselines
-rDPO (Chowdhury, Kini, and Natarajan 2024) does not use a DRO framework.
-Instead, it assumes that the observed preference labels are flipped at a known rate $\epsilon\in[0,0.5)$.
-Under this assumption, they derive an unbiased estimator of the DPO loss by reweighting the terms as follows:
-$\frac{1}{1-2\epsilon}\left[(1-\epsilon)\ell(x,y_{1},y_{2}1)-\epsilon\ell(x,y_{1},y_{2},-1)\right],$
-where $y_{1}$ is the chosen response in the noisy data.
-DrDPO applies DRO over the empirical data distribution using a KL-divergence-based ambiguity set:
-$\min_{\pi_{\theta}}\max_{P\in\mathcal{Q}(\mathcal{D})}\mathbb{E}_{(x,y_{1},y_{2},c)\sim P}[\mathcal{L}_{\text{DPO}}(\pi_{\theta})],$ (5)
-where $\mathcal{Q}(\mathcal{D}):=\left\{P:D_{\text{KL}}(P\|\mathcal{D})\leq\rho\right\}$. Using duality, the inner maximum (the objective to minimize) is approximated by:
-$\beta^{\prime}\log\mathbb{E}_{(x,y_{1},y_{2},c)\sim\mathcal{D}}\left[\exp\left(\frac{1}{\beta^{\prime}}\mathcal{L}_{\text{DPO}}(\pi_{\theta})\right)\right],$ (6)
-where $\beta^{\prime}$ is a hyperparameter.
-E.2 Whittle Index Policy
-Whittle index is a heuristic that is asymptotically optimal for planning in Multi-Armed Restless Bandits. To calculate whittle index, first we define the value function as
-$V_{i}^{\lambda}(s)=\max_{a\in\{0,1\}}Q_{i}(s,a_{i},\lambda).$ (7)
-Here, $Q_{i}(s,a_{i},\lambda)$ measures the expected discounted cumulative future reward where a reward compensation $\lambda$ is added to the reward when the passive action is taken.
-The Whittle index associated to the state $s_{i}$ is then defined as:
-$W_{i}(s_{i}):=\inf_{m}\left\{Q_{i}(s_{i},a_{i}=0,m)=Q_{i}(s_{i},a_{i}=1,m)\right\}.$
-The Whittle Index intuitively captures the value of taking an active action on an arm. We compute the Whittle Index using a binary search algorithm proposed by Qian et al. 2016.
-Finally, as a policy, we choose at every timestep the beneficiaires with top-K whittle indices.
-E.3 ARMMAN Real World Data Attributes
-To conduct our experiments in the ARMMAN real-world setting, we use a dataset of 2100 beneficiaires as a subset from a quality improvement study conducted in January 2022.
-Specifically, we use transition probabilities of beneficiaries estimated from historical data and use these probabilities to simulate beneficiary behaviour.
-Additionally, for every beneficiaries, we have information in their sociodemographic attributes, registration information, as well as preference on time slots of calls and on the language in which they want to receive calls.
-Table 5 shows the beneficiary attributes we use as the set of features that LLM can use in the reward function.
-Table 5: Features describing beneficiaries in the ARMMAN dataset
-Feature Name Data Type
-Enrollment gestational age Int
-Enrollment delivery status Int
-Gravidity (number of pregnancies) Int
-Parity (number of viable pregnancies) Int
-Live births count Int
-Days to the first call Int
-Ages 10–20 Binary
-Ages 21–30 Binary
-Ages 31–40 Binary
-Ages 41–50 Binary
-Ages 51–60 Binary
-Speaks Hindi Binary
-Speaks Marathi Binary
-Speaks Gujurati Binary
-Speaks Kannada Binary
-Education level 1/7 – illiterate Binary
-Education level 2/7 – 1–5th Grade Completed Binary
-Education level 3/7 – 6–9th Grade Completed Binary
-Education level 4/7 – 10th Grade Passed Binary
-Education level 5/7 – 12th Grade Passed Binary
-Education level 6/7 – Graduate Binary
-Education level 7/7 – Post graduate Binary
-Phone owner 0 (e.g., woman) Binary
-Phone owner 1 (e.g., husband) Binary
-Phone owner 2 (e.g., family) Binary
-To be called from 8:30am–10:30am Binary
-To be called from 10:30am–12:30pm Binary
-To be called from 12:30pm–3:30pm Binary
-To be called from 3:30pm–5:30pm Binary
-To be called from 5:30pm–7:30pm Binary
-To be called from 7:30pm–9:30pm Binary
-NGO Binary
-ARMMAN Binary
-PHC Binary
-Income bracket -1 (no income) Binary
-Income bracket 1 (0–5000) Binary
-Income bracket 2 (5001–10000) Binary
-Income bracket 3 (10001–15000) Binary
-Income bracket 4 (15001–20000) Binary
-Income bracket 5 (20001–25000) Binary
-Income bracket 6 (25001–30000) Binary
-Income bracket 7 (30000–999999) Binary
-E.4 RMAB Preference Data Generation
-To generate the preference dataset for fine-tuning an LLM, we follow these steps:
-  1. 1.
-    We consider a set of 190 human preference commands similar to the ones proposed in Behari et al. 2024. Table shows a smaller sample of those commands.
-  2. 2.
-    For each of these commands, we use the DLM pipeline to generate a set of 20 relevant reward functions using prompt in Figure 6.
-  3. 3.
-    We sample a set of 50 pairs from these reward functions. For each of these pairs
-    1. (a)
-      solve the RMAB problem and generate trajectories for every reward function
-    2. (b)
-      ask an LLM Judge to decide the better reward function using the prompt in Figure 7.
-  4. 4.
-    We perform this query 10 times to estimate LLM’s uncertainty in preference for every pair of chosen and rejected reward functions.
-E.5 Hyperparameters
-Table 6 shows the hyperparameters used during training.
-Table 6: Hyperparameter Configuration
-Description RMAB UltraFeedback
-Model Configuration
-Maximum sequence length 2048 1536
-LoRA rank parameter 64 64
-LoRA scaling parameter 16 16
-LoRA dropout rate 0.05 0.05
-Data type bf16 bf16
-SFT Training
-Training epochs 2 2
-learning rate 1e-5 1e-6
-per device train batch size 8 4
-number of devices 4 4
-gradient accumulation steps 1 2
-optimizer adamw_torch adamw_torch
-weight decay 0.01 0.01
-optimizer warmup steps 100 100
-learning rate scheduler cosine cosine
-DPO Training
-training epochs 1 1
-learning rate 1e-5 1e-6
-per device train batch size 4 2
-gradient accumulation steps 4 4
-optimizer warmup steps 100 100
-beta parameter 0.25 0.25
-optimizer adamw_torch adamw_torch
-warmup steps 100 100
-learning rate scheduler cosine w/min_lr=1e-6 cosine w/min_lr=1e-7
-Appendix F Additional Results and Ablations
-Figure 5: RMAB environment win rates under adversarial label noise
-F.1 Adversarial Noise
-In the main text we defined and presented results for label flipping noise. Here we consider an additional noise distribution defined as
-[…]
-$U$ $\sim\mathrm{Unif}(0,\alpha)$
-Similar to label flipping, adversarial noise tends to flip the labels, however, the mechanism is different as the distribution shift occurs as continuous change with an $\alpha$-ball of $q$.
-The results of implementing DPO-PRO and the baselines with adversarial noise in the UltraFeedback dataset are shown in Table 7(b), consistent with the trends we observed with label flipping where additional noise improves training in general, and DPO-PRO being the best performer.
-Figure 5 shows the results for the public health (RMAB) application.
-We show the former results in a table and the latter in a table for consistency with the main text.
-Table 7: Results on UltraFeedback under label-flip noise. Experiments run with Phi-3-mini. Mean and standard error over 2000 evaluation samples are shown.
-(a) Win rate ($\uparrow$)
-Method No noise Low noise (0.1) High noise (0.3)
-[…]
-[…]
-[…]
-[…]
-(b) Generation Evaluation Reward ($\uparrow$)
-Method No noise Low noise (0.1) High noise (0.3)
-[…]
-[…]
-[…]
-[…]
-F.2 Additional Ablations
-Effect of no Robustness ($\rho=0$)
-Table 8 presents results for a variant of DPO-PRO with $\rho=0$, which effectively removes the DRO component.
-In this case, the method reduces to a form of DPO that uses soft preference probabilities $q$ instead of binary labels.
-Interestingly, the performance remains slightly better than vanilla DPO.
-We note that when $\rho=0$, the loss becomes a standard cross-entropy between probability scores.
-Table 8: Flip Noise - Win Rate (%) and Evaluation Reward Comparison. DPO-PRO with $\rho=0$.
-Method No noise $\uparrow$ Low noise (0.1) $\uparrow$ High noise (0.3) $\uparrow$
-Win Rate (%) Eval Reward Win Rate (%) Eval Reward Win Rate (%) Eval Reward
-[…]
-[…]
-[…]
-UltraFeedback results on Llama3 8B
-Table 9(b) shows the results of fitting the LLama 3 8B model instead of the Phi3-mini 3B model.
-As expected the results are slightly higher considering the larger and more capable initial neural network.
-However, the memory and computation requirements also increase.
-Table 9: Results on UltraFeedback under label-flip noise run with Llama3 8B. Mean and standard error over 2000 evaluation samples are shown.
-(a) Win rate ($\uparrow$)
-Method No noise Low noise (0.1) High noise (0.3)
-[…]
-[…]
-[…]
-[…]
-(b) Generation Evaluation Reward ($\uparrow$)
-Method No noise Low noise (0.1) High noise (0.3)
-[…]
-[…]
-[…]
-[…]
-Appendix G Prompts
-table]box:rwgeneratorLLM prompt to generate reward function Prompt
-Create a Python reward function for RL in phone call resource allocation to mothers in India, with the objective of prioritizing higher states and: {goal_prompt}.
-The function should use ’state’ (value is either 0 or 1) and features ’agent_feats’ to direct the RL agent.
-Here is a description of the features you may use along with the index in the ’agent_feats’ array: {PROMPT_FEAT_DESC_DICT}
-Your task:
-1. Write a simple, single-line Python reward function. - Exclude the word ’return’. - Exclude non-standard libraries. - Format your code with []: [YOUR FUNCTION].
-2. Note that HIGHER states are always preferred, so ensure the reward increases as ’state’ increases.
-3. Make sure the reward is always positive and increasing with state.
-4. Avoid using bitwise operators. Use ’and’, ’or’ instead.
-Here is an example to guide you: PROMPT_REW_EXAMPLE_DICT
-Come up with a unique new reward for the specified goal: goal_prompt.
-Figure 6: Prompt passed to the LLM to generate a reward function based on the context of problem scenario in Real World Domain.
-table]box:rwgeneratorLLM Judge to Decide Preferred Response Prompt
-My goal was to create a Python reward function for RL in resource allocation, with the objective of: Focus on those with high education.
-I tried several reward functions for this task.
-Below, I have the given reward function, and the corresponding distribution of reward achieved.
-Below are the reward functions I used and their corresponding reward distributions: Function Number 0: Reward Function: -agent_feats[5] -agent_feats[6]-agent_feats[7]-agent_feats[8]-agent_feats[9]-agent_feats[10]-agent_feats[11]
-Reflection: ’
-Category: Age Ages 10-20: 121.73
-Ages 21-30: 421.04
-Ages 31-40: 244.49
-Ages 41-50: 64.11
-Ages 51-60: 10.58
-Category: Income Income bracket 1 (e.g., 0-5000): 126.82
-Income bracket 2 (e.g., 5001-10000): 373.62
-Income bracket 3 (e.g., 10001-15000): 234.87
-Income bracket 4 (e.g., 15001-20000): 77.40
-Income bracket 5 (e.g., 20001-25000): 35.58
-Income bracket 6 (e.g., 25001-30000): 2.58
-Income bracket 7 (e.g., 30000-999999): 11.09
-Category: Education Illiterate: 39.91
-1-5th Grade Completed: 157.84
-6-9th Grade Completed: 281.36
-10th Grade Passed: 197.64
-12th Grade Passed: 103.18
-Graduate: 21.13
-Post graduate: 60.89
-Function Number 1: Reward Function: state * agent_feats[10]
-Reflection: ’ Category: Age Ages 10-20: 134.22
-Ages 21-30: 469.16
-Ages 31-40: 270.44
-Ages 41-50: 72.80
-Ages 51-60: 11.96
-Category: Income Income bracket 1 (e.g., 0-5000): 138.40
-Income bracket 2 (e.g., 5001-10000): 414.44
-Income bracket 3 (e.g., 10001-15000): 266.44
-Income bracket 4 (e.g., 15001-20000): 85.33
-Income bracket 5 (e.g., 20001-25000): 40.20
-Income bracket 6 (e.g., 25001-30000): 2.80
-Income bracket 7 (e.g., 30000-999999): 10.96
-Category: Education Illiterate: 45.07
-1-5th Grade Completed: 173.82
-6-9th Grade Completed: 314.07
-10th Grade Passed: 217.31
-12th Grade Passed: 113.02
-Graduate: 29.36
-Post graduate: 65.93
-Based on the above reward distributions and the given goal: Focus on those with high education., please identify the FUNCTION NUMBER of the most effective reward function.
-Provide your answer EXACTLY IN the following format: ’The best reward function is at number: [FUNCTION NUMBER]’.
-Output:
-The best reward function is at number: 1
-Figure 7: Prompt passed to the LLM to choose a reward function based on the context of problem scenario in Real World Domain, the generated reward functions and the reward distribution corresponding to every reward function.
+Acknowledgements
+This work was supported by ONR MURI N00014- 24-1-2742 and based upon work supported by the AI Research Institutes Program funded by the National Science Foundation under the AI Institute for Societal Decision Making (NSF AI-SDM), Award No.
+2229881.